### v1

# Tumor-anchored deep feature random forests for out-of-distribution detection in lung cancer segmentation

Abstract
Accurate segmentation of cancerous lesions from 3D computed tomography (CT) scans is essential for automated treatment planning and response assessment. However, even state-of-the-art models combining self-supervised learning (SSL) pretrained transformers with convolutional decoders are susceptible to out-of-distribution (OOD) inputs, generating confidently incorrect tumor segmentations, posing risks for safe clinical deployment. Existing logit-based methods suffer from task-specific model biases, while architectural enhancements to explicitly detect OOD increase parameters and computational costs. Hence, we introduce a plug-and-play and lightweight post-hoc random forests-based OOD detection framework called RF-Deep that leverages deep features with limited outlier exposure. RF-Deep enhances generalization to imaging variations by repurposing the hierarchical features from the pretrained-then-finetuned backbone encoder, providing task-relevant OOD detection by extracting the features from multiple regions of interest anchored to the predicted tumor segmentations. Hence, it scales to images of varying fields-of-view. We compared RF-Deep against existing OOD detection methods using 1,916 CT scans across near-OOD (pulmonary embolism, negative COVID-19) and far-OOD (kidney cancer, healthy pancreas) datasets. RF-Deep achieved AUROC >> 93.50 for the challenging near-OOD datasets and near-perfect detection (AUROC >> 99.00) for the far-OOD datasets, substantially outperforming logit-based and radiomics approaches. RF-Deep maintained similar performance consistency across networks of different depths and pretraining strategies, demonstrating its effectiveness as a lightweight, architecture-agnostic approach to enhance the reliability of tumor segmentation from CT volumes.
keywords:
Tumor segmentation, lung cancer, computed tomography, out-of-distribution detection, random forests, outlier exposure
journal: Medical Image Analysis
[…]
organization=Department of Medical Physics
Memorial Sloan Kettering Cancer Center,addressline=1275 York Avenue, city=New York City, postcode=10065, state=New York, country=USA
1 Introduction
Deep learning (DL) based medical image analysis is now applicable throughout the continuum starting from diagnosis, treatment planning, and response monitoring at follow-up to enhance care of patients with cancer. For instance, hybrid transformer-convolutional networks have shown to provide highly accurate tumor and healthy tissue segmentations from radiographic modalities (jiang2022self; Willemink2022; NguyenAAAI2023; YanWACV2023; Qayyum2023BHI; gu2024build). DL methods for medical applications are often developed using narrowly scoped datasets. Consequently, accuracy on in-distribution (ID) testing doesn’t translate to out-of-distribution (OOD) clinical cases impacted by different imaging acquisitions (Roschewitz2023Automatic; Koch2024), concept drifts (huang2022developing; sahiner2023data; gomez2024swintransformers), and label shifts like varying delineation guidelines (Nichyporuk2022MELBA; Lempart2023JACMP; Godau2025). Importantly, DL models also produce confidently inaccurate predictions (hendrycks2018deep; quinonero2022dataset; prabhu2022adapting_Nips2022; banerjee2023shortcuts), hindering ability of clinicians to leverage model uncertainties in utilizing AI predictions for clinical decision making (otles2021shift). Over-reliance on AI models by experts carries the risk of clinical expert deskilling (Budzyn2025Deskilling), exacerbating risk of harm to patients. These considerations underscore the critical need to provide reliably accurate methods that can contextualize task model predictions with respect to data instances used in prediction.
Segmentation accuracy metrics like Dice and Hausdorff distances, though useful to compare models’ performance on data cohorts, do not reflect well-calibrated uncertainty estimates even on ID datasets (Ren2024; Rangnekar2025SPIE). Model uncertainty visualization (goddard2012automation) has been shown to enhance consistency of clinical delineations (DeBiase2024). However, such measures can be unreliable in OOD scenarios, as demonstrated in three out of four OOD cases in Figure 1. Methods to explicitly flag OOD scans are often narrowly scoped to detect any variation from training data (Roschewitz2023Automatic; Godau2025); the model itself cannot generalize to common image variations not encountered in training, limiting their use for practical purposes. Our work overcomes the limitations of prior approaches by leveraging pretrained models to ensure sufficient generalization to common clinical imaging variations while enhancing robustness to far-OOD (disease in different anatomical sites) and near-OOD (different disease in same anatomic site) use cases. Our contributions are:
Figure 1: Uncertainty maps for in-distribution (ID) and out-of-distribution (OOD) scans. (a) ID lung cancer cases (1–2) show concentrated boundary uncertainty; OOD cases (3–6: pulmonary embolism, COVID-19-negative, kidney cancer, healthy abdomen) exhibit concentrated, diffused, or misaligned patterns.
  • 1.
    A post-hoc OOD detection approach using random forests with deep features (RF-Deep), trained with outlier exposure. Features are extracted from regions of interest (ROIs) anchored to the task-specific model’s predictions, enabling scan-level and task-relevant OOD filtering. Our approach reduced false detections, achieving AUROC >> 90 on challenging near-OOD datasets and near-perfect detection on the far-OOD datasets.
  • 2.
    Generalizability to common imaging variations is ensured by using features extracted from the hierarchical encoder layers of a segmentation model subject to self-supervised learning (SSL) pretraining and task-specific fine-tuning. Robustness of task-specific model performance on commonly encountered imaging variations is also reported.
  • 3.
    Evaluation on 2,056 3D CT scans from five public datasets, with analysis of SSL pretraining methods, model depths, and classifier designs on OOD detection performance.
  • 4.
    Analysis of why RF-Deep outperforms radiomics-based and logit-based approaches by performing unsupervised clustering, SHAP-based feature importance, and spatial uncertainty analysis.
2 Related Works
Broadly, there are two classes of approaches for improving the generalizability of models for OOD detection. The first class consists of learner-based methods that leverage large corpus of training data, regularization techniques, and specialized network architectures to extract feature representations that are robust to data variations. The second class are filter-based methods that explicitly flag OOD cases at test time, preventing unreliable predictions and thereby improving overall reliability.
2.1 Learner-based methods
Learner-based methods focus on extracting robust feature representations from available training data such that the model can be applied across different domains without performance degradation. Domain-invariant feature representations can be extracted using: (a) specialized architectures (Saha2021EndToEndProstate; Simeth2023DominantIndexLesion), (b) data augmentation methods in image and feature space (Jung2019ProstateMixup; Chen2022EnhancingMRImageSegmentation; vaish2025dataagnostic), (c) regularized training with multi-task objectives (Panfilov2019RobustnessKneeMRI; Wen2024FromDenoisingTraining), as well as (d) SSL pretraining with numerous unlabeled examples to expose models to wider ranges of imaging and data variations than is practical for supervised training (hatamizadehSwinUNETR2022; haghighi2022dira; jiang2024self; Du2024SegVol).
Prior works on tumor segmentation have shown that use of enhanced network architectures combining attention mechanisms and residual connections that effectively increase network size produce robust generalization across different MR acquisitions (Saha2021EndToEndProstate; Simeth2023DominantIndexLesion). Another approach by Gu2021_DCANet_MICCAI explicitly trained a model to extract separate basis functions for various domains that were then combined using group convolutions to extract a generalizable representation applicable to unseen image domains for prostate gland segmentation.
An alternative approach to creating specialized architectures is to employ data augmentation and regularization tasks when training standard networks like the U-Net (Chen2022EnhancingMRImageSegmentation; Wen2024FromDenoisingTraining). Mixup is a type of data-agnostic augmentation strategy applicable either to the inputs, labels, as well as feature representations extracted from the input images (Jung2019ProstateMixup; vaish2025dataagnostic). Advanced feature augmentation strategies using normalizing flows extract feature statistics for creating augmented features in individual layers, enhancing domain generalization for prostate gland and eye fundus segmentation (SafRez_MixStyleFlow_MICCAI2025). Conceptually, the mixup and augmentation methods are general and could be combined with transformers to further enhance robustness to distribution shifts (Boone2023ROODMRI).
SSL pretraining is an altogether different approach compared to prior methods that employs multi-stage training to fully exploit the image information to extract distribution shift robust feature representations. This is accomplished either through pretraining with extensive data augmentations (Fedorov2021Tasting) or simply by exposing the model to a large corpus of unlabeled examples encompassing possible imaging variations (hatamizadehSwinUNETR2022; jiang2024self; gomez2024swintransformers). A different approach is a successive training strategy that explicitly addresses generalization to OOD by converting a portion of every encountered OOD data to ID data and successively refines the model to that dataset (Azizi2023REMEDIS). Unlike SSL, which aims to reduce need for large downstream datasets, this requires data for adapting to every new domain.
2.2 Filter-based OOD methods
Whereas learner-based methods seek to enhance models’ consistency across domains subject to distribution shifts, filter-based methods aim to enhance the confidence of predictions by filtering out data that differs from the training distribution. For this purpose, filter-based methods implement either a separate OOD model or a task-integrated OOD model to identify OOD data scans. Scans detected as OOD are removed, ensuring that models are applied only on ID-like data where predictions and uncertainties are valid, providing reliably safe performance required in clinical settings (berger2021confidence).
Similar to learner-based methods, filter-based methods have been applied to handle imaging acquisition differences (tardy2019uncertainty; gao2020response; nandy2021distributional). The vast majority of filter-based methods focused on disease classification from 2D images; method complexity varied from easy-to-compute Mahalanobis distances (tardy2019uncertainty), dedicated networks for detecting OOD cases (nandy2021distributional) to more complex instance-level retraining performed to assign metric distance to OOD scans by measuring response variations on the ID test set (gao2020response). A different approach involves using a single model to classify tasks on inlier data, while exposing the same model to outlier examples to detect OOD during inference (araujo2023few; roy2022does). The last approach is practical to implement when OOD cases are known as domain knowledge (e.g. optical tomography versus retinal fundus images) (araujo2023few), or when the outliers like infrequent disease conditions are deducible from the training as OOD (roy2022does).
Similar to classification tasks, Mahalanobis distance computed from feature embeddings has been used to detect OOD for lung lesion segmentation on CT scans of patients with COVID-19 (gonzalez2022distance). However, the high dimensionality of feature embeddings limits accuracy (roy2022does), due to difficulty in distinguishing subtle differences in appearance of different pathologies occurring in the same disease site (berger2021confidence). Dimensionality reduction methods including PCA (woodland2024dimensionality) as well as VQ-GAN (graham2023unsupervised; pinaya2022unsupervised) have been used to reduce the latent space and sharpen the separation of ID from OOD feature distributions. However, dimensionality reduction can suffer from high inductive bias (nalisnickMTGL19), reducing applicability to wider range of OOD scans. These dimensionality reduction approaches, including GANs, also require additional training with secondary datasets. Calibrated uncertainty estimation methods using computationally simple temperature scaling (karimi2022improving) as well as logit-based OOD scores (hendrycks17baseline; hendrycks2022logits) have been used for OOD detection. However, effectiveness of such methods are limited by inherent bias of the task-specific model use to estimate uncertainty (mehrtash2020confidence). Modified architectural formulations (yuan2023devil; Larrazabal2023MEEP) and generative diffusion methods (Nguyen2026) to detect OOD increase network parameters and require additional training. Finally, radiomic feature distance methods though useful for OOD detection at cohort level (konz2024radmetricmedicalimage; vasiliuk2023limitations) are not applicable to scan-level detections due to their difficulty in extracting subtle data variations while providing robust performance over common variations.
3 Framework overview and OOD task definition
Our goal for OOD detection was to develop a highly accurate scan-level OOD detection approach, which does not require large numbers of outlier samples, with generalizable applicability to images with different acquisitions and fields of view. Let $\mathcal{D}_{\text{in}}$ denote the distribution of lung cancer CT scans, closely matching the training dataset used to create the lung tumor segmentation model, and let $\mathcal{D}_{\text{out}}$ represent the scans that differ in pathology and anatomic sites. The objective is to determine whether a new scan $x$ belongs to $\mathcal{D}_{\text{in}}$ or $\mathcal{D}_{\text{out}}$ at the scan level via a scoring mechanism provided by the random forest classifier.
Our approach combines the advantages of learner-based methods (by using the backbone of a segmentation model subject to SSL pretraining to extract imaging domain-agnostic features) with a filter-based classifier to reject conceptually different scans. An architecturally simple and lightweight random forest OOD classifier incorporates known outliers during training (hendrycks2018deep; thulasidasan2021effective; GuhaRoy2022) to enhance the classifier’s ability to flag unseen outliers. Scan-level OOD prediction for segmentation involves thousands of voxel-level predictions, which, while offering rich spatial context, also introduces memory constraints in processing and aggregating information from their large 3D volumes. Our approach addresses this issue by focusing the OOD detection in regions of interest (ROI) anchored to the predicted tumor segmentation, and then averaging the metric scores from those regions, which also allowed the approach to scale to images with varying field of views. Using regions anchored near the predicted tumors also allows the model to focus on differentiating features in relevant parts of the image, as opposed to distinguishing entire images, which in turn allows it to distinguish pathologies occurring in the same disease site.
3.1 Segmentation model
The segmentation framework employs a hybrid transformer-convolution architecture to leverage strengths of global contextual modeling with the transformer backbone as encoder and local spatial precision of the convolutional decoder. The backbone is a hierarchical Swin Transformer (liu2021swin) with depth configuration 221222-2-12-2 across four successive stages, with a patch size of $2\times 2\times 2$ voxels and a window size of $4\times 4\times 4$ voxels. Its hierarchical design progressively aggregates features at multiple spatial scales to capture fine-grained anatomical details. Windowed self-attention within each stage reduces complexity while preserving details and global context, critical for processing large volumetric images of size $128\times 128\times 128$ voxels. The backbone was pretrained with the self-distilled masked image transformer (SMIT) framework to enhance robustness to imaging and patient variations (jiang2022self). SMIT combines masked image prediction (he2021masked; xie2022simmim) with self-distillation (zhou2021ibot); described in Appendix A. The decoder, a 3D U-Net-based (ronneberger2015u) convolutional structure, was randomly initialized and fine-tuned with ID task-specific dataset. Following training, the entire model is frozen ([Uncaptioned image]) for subsequent segmentation and features-based OOD detection.
Figure 2: RF-Deep workflow for scan-level OOD detection. Panels (a–c) depict feature extraction from the frozen Refer to caption segmentation model and random forest training in an outlier exposure manner and panel (d) shows scan-level ID/OOD inference using the trained detector.
3.2 Scan-level OOD detection approach
Ou OOD detection pipeline consists of the following four steps (Figure 2):
  • 1.
    Regions of interest extraction: The task-specific segmentation model (Subsection 3.1) was used to generate segmentation of tumors from the entire scan. Detected tumor regions from each scan were used to extract multiple 3D ROIs, ensuring the tumor appears within each ROI volume without precise centering (Figure 2a).
  • 2.
    Feature extraction. We repurposed the fine-tuned backbone encoder from the segmentation model to extract multi-scale features from the five encoder stages: the patch embedding layer (48 channels) and the four Swin Transformer stages (96, 192, 384, and 768 channels, respectively). All features were aggregated into a single scan-level descriptor through global average pooling (GAP) describing the scan as a vector of features (Figure 2b).
  • 3.
    Training the OOD detector. Feature vectors were extracted from ID lung cancer datasets as well as a representative set of OOD scans from tumor anchored regions as described above and used to train the RF classifier for each cohort once. This approach enabled our detector to learn subtle decision boundaries that distinguish shifts even between different diseases affecting the same anatomic region (Figure 2c).
  • 4.
    Online inference. Given a new test instance, the aforementioned pipeline extracts potential tumor regions and tumor-anchored ROIs, followed by the extraction of GAP features, which the RF classifier uses to generate scan-level OOD classification (Figure 2d).
4 Experiments setup
4.1 Implementation details
All models were implemented in PyTorch (paszke2019pytorch) and MONAI (cardoso2022monai). All scans were intensity-normalized using the lung window (Hounsfield units (HU): [400,400][-400,400]) and resampled to a uniform voxel spacing of $1~\text{mm}^{3}$.
Segmentation. Models were fine-tuned using cross-entropy and Dice loss with a batch size of 16 distributed across four NVIDIA GPUs, with an input size of 128 $\times$ 128 $\times$ 128 voxels. We used a learning rate of $2\times 10^{-4}$ with linear warm-up (50 epochs) and cosine annealing over 1,000 epochs. Data augmentations included random flips, rotations, affine transformations, and illumination adjustments, consistent with tang2022self; jiang2022self. For inference on the full scan volume, we used a sliding window with 50% Gaussian overlap to obtain the segmentations.
OOD Detection. Random forest classifiers employed 1,000 trees with a maximum depth of 20 and balanced class weights (using default scikit-learn hyperparameters). We extracted deep feature representations from $n=4$ tumor-anchored 3D ROIs per scan for all ID and OOD datasets. All ROIs were used individually during training to increase training samples, and probabilities from all ROIs were averaged at the scan level as an ensemble for inference. For MaxSoftmax, MaxLogit, and energy, scan-level scores were obtained by mean-aggregating tumor (positive class) predictions across all voxels using the corresponding metric.
4.2 Datasets
Backbone pretraining. The backbone encoder was pretrained using the self-distilled masked image transformer (SMIT) framework on 10,432 3D CT scans, encompassing public and institutional datasets, following jiang2022self; tang2022self; jiang2024self.
Comparison to other pretraining strategies and architectures. Additionally, we pretrained using SimMIM (xie2022simmim), iBOT (zhou2021ibot), and Swin UNETR (tang2022self) to assess the impact of pretraining strategy on both accuracy and OOD robustness. To analyze the effect of model depth, we pretrained a lite backbone configuration (22222-2-2-2) using SMIT. Finally, we evaluated the publicly available Swin UNETR pretrained checkpoint (marked as ) to assess the impact of pretraining data.
Fine-tuning. We combined the pretrained backbone as the encoder with a randomly initialized U-Net decoder and fine-tuned using $N=317$ 3D scans from publicly available non-small cell lung cancer (NSCLC) Radiomics dataset (aerts2015data) with radiologist-provided tumor delineations.
Testing. We used the NSCLC Radiogenomics dataset (bakr2017data), consisting of 140 images from various scanners (GE, Siemens), as the ID test set. OOD evaluation used 1,916 3D CT scans from four public datasets: pulmonary embolism (RSNA PE, N=1,225N=1,225(colak2021rsna), COVID-19 negative patients (MIDRC C19, N=120N=120(ricord2021covid), kidney cancer (KiTS, N=489N=489(heller2023kits21), and healthy scans of the pancreas (PancreasCT, N=82N=82(roth2015deeporgan). The near-OOD datasets (RSNA PE, MIDRC C19) consist of chest CTs involving non-cancerous pathologies such as pulmonary embolism and inflammation, depicting appearances similar to lung lesions. The far-OOD datasets (KiTS, PancreasCT) include cases with larger field-of-view than training with healthy organs and non-lung cancers.
4.3 Evaluation metrics
Segmentation accuracy on ID datasets was measured using the Dice similarity coefficient (DSC) and the 95th percentile Hausdorff distance (HD95). OOD detection accuracy was computed using the area under the receiver operating characteristic curve (AUROC) and false positive rate at 95% true positive rate (FPR at 95%TPR, or FPR95). AUROC quantifies the separability between ID and OOD cases (davis2006relationship), with higher values indicating better discrimination and FPR95 measures the proportion of OOD cases misclassified as ID when sensitivity is fixed at 95% (liang2017enhancing), where lower values are better.
To ensure meaningful AUROC computation, the number of OOD scans was balanced with ID test cases ($N=140$) by sampling OOD examples with replacement in datasets (RSNA PE and KiTS) with a much larger number of OOD scans and averaging the computed metrics across multiple draws. Splits ensured zero patient overlap with consistent OOD case counts across runs. This evaluation approach avoids biases from imbalanced evaluation sets while ensuring stable ID-OOD ratios, following best practices for robust OOD evaluation (szyc2023out; humblot2023beyond).
For segmentation performance, we reported the mean and standard deviation across test cases. For OOD detection performance, we reported the point estimates with 95% bootstrap confidence intervals computed over 100 runs with matched random seeds to reflect sampling variability across different train-validation splits.
4.4 Comparative OOD detection methods
Baseline comparisons included commonly used OOD detection methods, namely, MaxSoftmax (hendrycks17baseline), MaxLogit (hendrycks2022logits), and energy (liu2020energy), performed by computing scan-level scores by mean-aggregating the tumor (positive class) predictions across all voxels (see Appendix B).
In addition, a RF classifier was created using handcrafted radiomics features ($N=293$) called RF-Radiomics, computed within the segmented tumor regions following the Image Biomarker Standardization Initiative (IBSI) guidelines (zwanenburg2020image) with open-source software PyCERR (apte2018extension). Details of individual radiomic features are in Table 5. The high correlation among radiomic features was addressed by employing recursive feature elimination to select a more compact and informative set of radiomic features. Of note, recursive feature elimination was not used for training the RF-Deep classifier.
Finally, the Mahalanobis distance-based OOD detector (MD-Deep) (lee2018simple) was computed with respect to deep features of the ID data, modeled as a single Gaussian in feature space using Ledoit-Wolf covariance estimation. Hence, MD-Deep computes OOD scores as the distance of the each scan with respect to the fitted ID distribution. We excluded ODIN (liang2017enhancing), as it requires temperature scaling and gradient perturbations on fine-tuning data, which in real clinical settings is often unavailable when applied to new OOD cohorts.
(a) Segmentation metrics
(b) Imaging variations
(c) Segmentation examples
Figure 3: Pretraining strategies performance and robustness evaluation on ID test set. (a) Segmentation metrics (DSC and HD95). (b) Performance across imaging variations. (c) Representative segmentations; R denotes reconstruction kernel.
5 Results
5.1 Segmentation performance on ID datasets
Segmentation performance for the various analyzed pretraining strategies is shown in Figure 3(a) for the ID testing dataset. All pretrained methods were similarly accurate, with the lowest accuracies produced by iBOT, followed by the publicly available Swin UNETR checkpoint.
Next, we performed a disaggregated analysis of model performance under common imaging variations including scanner manufacturer (GE versus non-GE), contrast versus non-contrast enhanced CT, and reconstruction kernels. Reconstruction kernels were categorized as Recon 1 (GE ‘standard’ and ‘bone’, Siemens with << B40), Recon 2 (GE ‘Bone Plus’, Siemens with $\geq$ B40 and << B50), and Recon 3 (GE ‘Lung’, Siemens $\geq$ B50). Model fine-tuning was performed on the public dataset acquired on Siemens scanners with contrast and Recon 1 kernel. From Figure 3(b), all models were similarly accurate across the analyzed variations, with larger gaps in accuracy noted for Swin UNETR ( Recon 1, non-contrast CT) and iBOT (Recon 3, non-contrast). All other strategies showed relatively small degradation in accuracy due to the analyzed variations indicating robust performance through any of the pretrained methods. SMIT and SMIT Lite strategies were similarly accurate, indicating that a smaller model is sufficient for robust tumor segmentation. On the other hand, choice of pretraining data impacted accuracy, as seen by the performance difference between the Swin UNETR trained with the 10K dataset and the public checkpoint Swin UNETR. Detailed accuracies using DSC and HD95 metrics for the analyzed methods are shown in Figure 10.
Overall, the SMIT-pretrained model was the most accurate, followed by the Swin UNETR method, and was thus selected for benchmarking the OOD detection method using RF-Deep througout this paper.
Table 1: OOD detection performance comparing various methods on the segmentation model fine-tuned using the SMIT-pretrained backbone. AUROC ($\uparrow$) and FPR95 ($\downarrow$) are reported with 95% bootstrap confidence intervals over 100 matched-seed runs.
Method RSNA PE MIDRC C19 KiTS PancreasCT
AUROC FPR95 % AUROC FPR95 % AUROC FPR95 % AUROC FPR95 %
MaxSoftmax
88.61
(84.62–92.36)
38.37
(31.78–49.24)
86.57
(81.79–90.14)
49.27
(39.53–59.78)
95.67
(93.41–98.06)
23.26
(11.47–34.19)
94.61
(92.61–97.49)
34.27
(24.80–49.30)
MaxLogit
88.77
(85.09–92.43)
40.01
(24.80–52.52)
89.31
(85.05–93.06)
41.59
(29.84–53.96)
95.89
(93.58–98.44)
18.05
(10.41–27.37)
93.53
(90.17–96.59)
30.87
(16.17–56.87)
Energy
88.59
(84.85–92.24)
39.86
(24.80–52.55)
89.30
(84.98–93.04)
43.04
(29.84–55.43)
95.80
(93.41–98.44)
17.47
(10.41–26.65)
93.52
(90.10–96.53)
29.96
(15.83–56.87)
RF-Radiomics
88.70
(81.50–93.00)
41.10
(25.20–57.80)
91.30
(87.10–94.40)
30.00
(17.40–45.60)
96.20
(93.00–98.30)
11.80
(5.80–21.90)
96.10
(89.10–99.20)
17.20
(4.30–41.90)
MD-Deep
87.30
(83.90–91.10)
45.10
(30.60–60.30)
84.50
(79.40–89.00)
58.50
(38.20–77.60)
99.20
(97.90–99.80)
3.10
(0.00–9.20)
99.20
(98.20–100.00)
2.80
(0.00–7.70)
RF-Deep
95.80
(92.60–98.50)
15.20
(7.10–23.50)
93.30
(89.90–96.00)
25.50
(16.80–36.20)
100.00
(99.90–100.00)
0.10
(0.00–1.00)
100.00
(100.00–100.00)
0.00
(0.00–0.00)
5.2 Out-of-distribution detection performance
RF-Deep is the most accurate method for both near-OOD and far-OOD datasets (Table 1). On the near-OOD datasets (RSNA PE and MIDRC C19), RF-Deep achieved AUROCs of 95.80 and 93.30, outperforming the next best method (MaxLogit) by approximately 7.7% and 4.0%, respectively. Similarly, RF-Deep reduced FPR95 by 18–25 absolute points compared to the next closest competitor. These improvements remain consistent within the 95% confidence intervals, with non-overlapping intervals confirming statistically significant superiority over all baseline methods, highlighting the stability of RF-Deep across bootstrap resampling.
RF-Deep also achieved near-perfect separability on the far-OOD datasets using KiTS (AUROC 100, FPR95 0.10%) for kidney cancers and PancreasCT (AUROC 100, FPR95 0.00%) containing healthy abdominal scans, effectively eliminating false alarms for anatomically distant pathologies. Although it was less accurate on the near-OOD cohorts compared to its accuracy on the far-OOD cohorts, the gap in AUROC (RSNA PE: ++3.5%, MIDRC C19: ++6.7%) was the smallest compared to all other methods, with the next smallest gap observed for RF-Radiomics (RSNA PE: ++7.5%, MIDRC C19: ++4.9%).
RF-Radiomics, while competitive on the far-OOD datasets (AUROC of 96.20 on KiTS and 96.10 on PancreasCT with relatively low FPR95), produced moderate error rates on the near-OOD cohorts (AUROCs 88.70 and 91.30, FPR95s around 3041%), indicating that handcrafted features are less effective than deep features at capturing subtle differences in the appearance of pathologies (analyzed further in Sections 5.3.1 and 5.3.2).
MD-Deep, although highly accurate for detection on far-OOD datasets (AUROC 99.2 on both KiTS and PancreasCT) was less effective and similar to logit-based methods on near-OOD cohorts (AUROC 87.3 on RSNA PE and 84.5 on MIDRC C19, with FPR95 around 45-59%). MD-Deep was also less accurate than RF-Deep on near-OOD datasets, despite both methods using identical encoder features, indicating that feature selection performed by the RF was essential to achieve better separation of ID from OOD scans. Additional evaluation to assess if feature reduction enhanced performance showed no improvement compared to the standard MD-Deep approach (Figure 11).
All logit-based methods showed a considerable drop in accuracy for the near-OOD datasets compared to their performance on far-OOD datasets, reflecting the challenge in detecting OOD from images depicting different pathologies but from the same disease site as the ID dataset. For instance, MaxLogit exhibited performance gaps of 7–8% between near-OOD and far-OOD datasets, more than twice the gap observed for RF-Deep (3.5–6.7%), suggesting that logit-based scores alone may be insufficient for challenging near-OOD detection. We analyzed the underlying causes of these performance differences in Section 5.3.3.
5.2.1 Impact of model pretraining on OOD detection
Next, we analyzed whether the choice of model pretraining, model size (SMIT versus SMIT Lite), and pretraining data (Swin UNETR versus Swin UNETR) impacted the accuracy of RF-Deep. RF-Deep remained the most accurate method across all backbone architectures (Table 6), consistently achieving near-perfect accuracy on the far-OOD datasets. Table 7 further demonstrated that RF-Deep’s superiority over logit-based and radiomics methods was consistent across all tested backbones, with RF-Deep outperforming the next best method by large margins on the near-OOD datasets. Accuracy differences between the SMIT and SMIT Lite models on the near-OOD cohorts were modest (RSNA PE AUROC: 93.7 vs 94.1; MIDRC C19 AUROC: 93.2 vs 90.4; RSNA PE FPR95: 17.6% vs 20.4%; MIDRC C19 FPR95: 26.5% vs 30.5%), with SMIT Lite showing slightly better performance on RSNA PE but marginally lower accuracy on MIDRC C19.
Similarly, RF-Deep performance for Swin UNETR and Swin UNETR backbones was comparable (RSNA PE AUROC: 90.4 vs 89.6; MIDRC C19 AUROC: 88.6 vs 88.9; RSNA PE FPR95: 41.5% vs 34.1%; MIDRC C19 FPR95: 35.8% vs 37.5%). However, both Swin UNETR methods exhibited 4–7% lower AUROC on the near-OOD datasets compared to SMIT-based models, reaffirming that pretraining strategy and encoder architecture substantially influence OOD detection performance within the RF-Deep framework.
5.3 Why RF-Deep outperforms alternatives
(a) Deep features
(b) Radiomics features
Figure 4: t-SNE projected embeddings showing dataset-wise separability of (a) deep features and (b) radiomics features. Results from one representative split (of 100), combining all datasets in a single visualization, are shown for brevity. The shaded blue regions indicate the convex hull of the ID dataset.
5.3.1 RF-Deep extracts features that are better separated than RF-Radiomics
Figure 4 visualizes the t-SNE embeddings (perplexity=30, 1,000 iterations) of deep features (panel a) and radiomics features (panel b) for the ID and OOD datasets collectively. RF-Deep features exhibit better cluster separation between the ID and OOD datasets, with minimal overlap between ID and anatomically distinct far-OOD datasets (KiTS, PancreasCT) and moderate overlap with near-OOD datasets (RSNA PE, MIDRC C19), consistent with their increased detection difficulty. Radiomics features produced poorly separated clusters for the various datasets particularly for the near-OOD datasets with respect to the ID cases, explaining the relatively lower accuracies (Table 1), and less compact OOD score distributions than RF-Deep between ID and OOD datasets as shown in Figure 5.
(a) Radiomics features
(b) Deep features
Figure 5: Scan-level OOD detection score distributions for RF-Radiomics (a) and RF-Deep (b) across 100 matched-seed runs.
We next evaluated whether deep features produced by other pretraining strategies achieved similar separability as the SMIT-pretrained model. As shown in Figure 12, all pretraining methods improved separability compared to handcrafted radiomics features. However, class separation was slightly better using SMIT-based RF-Deep compared to SMIT Lite-based RF-Deep. The differences in class separation were more pronounced between Swin UNETR and Swin UNETR, than among SMIT, SimMIM, and iBOT, indicating the larger impact of pretraining data rather than the pretraining approach itself.
5.3.2 RF-Deep extracts OOD data-agnostic and specific features for improved separability
We used SHAP (SHapley Additive exPlanations) (lundberg2017unified) to identify which features drive RF-Deep and RF-Radiomics predictions, revealing how the two approaches differ in their decision-making mechanisms. Deep features (Figure 6) show broad, dataset-specific spread of important dimensions, with some features reused across multiple datasets (Ft 65, Ft 131, Ft 179, Ft 319), suggesting they capture general distributional properties relevant to distinguishing ID from OOD scans. Other features activate selectively for specific distribution shifts (e.g., Ft 5 and Ft 1381 for far-OOD), indicating that RF-Deep adapts its decision boundaries to the characteristics of each OOD scenario rather than relying on a fixed, universal subset of features. Notably, only a small fraction of the 1,488 encoder features appear consistently useful across OOD datasets, underscoring the need for outlier data exposure, especially in near-OOD scenarios where subtle feature combinations are required to distinguish concept shifts.
By contrast, radiomics features (Figure 7) emphasized shape-based and texture-based descriptors, specifically the Gray-Level Co-occurrence Matrix (GLCM) and Gray-Level Run-Length Matrix (GLRLM) subsets. While radiomic features lend interpretability compared to deep features, they are insufficient to capture the subtle differences across the datasets to accurately detect OOD.
(a) RSNA PE
(b) MIDRC C19
(c) KiTS
(d) PancreasCT
Figure 6: SHAP feature importance for RF-Deep across OOD datasets using the SMIT-pretrained segmentation model. Different deep features are activated for different distribution shifts.
(a) RSNA PE
(b) MIDRC C19
(c) KiTS
(d) PancreasCT
Figure 7: SHAP feature importance for RF-Radiomics across OOD datasets using segmentations from the SMIT-pretrained model. RF-Radiomics primarily relies on GLCM and GLRLM texture descriptors.
5.3.3 Why logit-based spatial uncertainties are less accurate than RF-Deep?
Figure 8: Representative MaxLogit visualizations for (a) segmentation predictions, (b) spatial heatmaps, and (c) boundary vs. interior distributions with predicted tumor volumes. Violet regions indicate unevaluated voxels due to foreground thresholding.
Table 2: MaxLogit scores across ID and OOD datasets within predicted tumor (positive class) regions, stratified into overall, boundary, and interior voxels.
Dataset Overall Boundary Interior
ID 2.63 $\pm$ 0.80 1.13 $\pm$ 0.35 3.25 $\pm$ 0.92
RSNA PE 1.42 $\pm$ 0.56 0.74 $\pm$ 0.25 1.92 $\pm$ 0.62
MIDRC C19 1.27 $\pm$ 0.70 0.82 $\pm$ 0.41 2.01 $\pm$ 0.73
KiTS 1.09 $\pm$ 0.40 0.85 $\pm$ 0.34 1.57 $\pm$ 0.49
PancreasCT 1.08 $\pm$ 0.60 0.66 $\pm$ 0.32 1.55 $\pm$ 0.75
Table 3: Classifier comparison on the pooled deep features. Point estimates of AUROC ($\uparrow$) and FPR95 ($\downarrow$) are reported over 100 matched-seed runs.
Dataset Linear Probing MLP Classifier Random forest
RSNA PE 92.76 / 28.00 93.01 / 19.31 95.80 / 15.20
MIDRC C19 92.36 / 29.46 92.88 / 27.88 93.30 / 25.50
KiTS 98.92 / 0.110 99.47 / 2.490 100.0 / 0.10
PancreasCT 99.94 / 0.460 99.25 / 0.800 100.0 / 0.00
To understand why logit-based OOD detectors were less accurate than RF-Deep, we analyzed MaxLogit spatial patterns within predicted tumor regions. We extracted boundary regions as a ring of 3 voxel thickness by performing erosion followed by dilation, each using a radius of 1 voxel. The interior contour following the erosion operation constituted the interior regions.
MaxLogit scores of ID and OOD scans show strong overall separation (Mann-Whitney U tests, all p << 0.001, effect sizes ranging from r = -0.75 for MIDRC C19 to r = -0.90 for KiTS) due to higher overall MaxLogit scores for ID cases than individual OOD datasets (Table 2). However, the scores have substantially overlapping ranges with ID (ID: 0.12–4.33, RSNA PE: 0.09–3.67, MIDRC C19: 0.30–2.70).
Concretely, interior voxels showed consistently higher MaxLogit scores than the boundary voxels in both ID and OOD predictions. For ID scans, the interior scores were roughly three times higher than boundary scores, and this ratio persists across all OOD datasets despite their lower absolute scores. Paired, two-sided Wilcoxon signed-rank tests comparing interior and boundary scores within each dataset confirmed the universality of this spatial pattern (p << 0.001 for all datasets, effect sizes r $\approx$ -1.0), indicating that the boundary–interior relationship is preserved regardless of ID or OOD status, further limiting their discriminative power.
Figure 8 provides a visualization of the MaxLogit spatial uncertainties for four representative cases. The density-normalized histograms depict the distribution of MaxLogit values in the interior and boundary regions of the detected lesions. The interior and boundary regions show very good separation for the ID case (row 1) and sufficiently different spatial distribution for the second case (row 2), enabling correct detection of OOD for this case. However, the third case (row 3) presents a challenge wherein the spatial uncertainty distribution closely resembles that of the ID case, resulting in incorrect classification. In contrast, RF-Deep correctly identified both OOD scans, demonstrating its superior ability to capture contextual features beyond local spatial patterns. Of note, both RF-Deep and MaxLogit-based OOD prediction was wrong for the fourth case (row 4), presenting a limiting case for either method.
An additional limitation of logit-based methods is their inherent scale invariance: MaxLogit scores reflect local appearance characteristics independent of lesion size as shown by the distributions for lesions of varying sizes in Figure 8. This is because MaxLogit computation treats voxels identically based on softmax probabilities, ignoring volumetric context that could differentiate true tumors from anatomically implausible detections. This scale invariance poses a critical bottleneck for OOD detection, as many false-positive detections on OOD scans manifest as small, scattered regions with locally high confidence but globally inconsistent spatial distributions. A method like RF-Deep on the other hand, aggregates features from regions encompassing detections, reducing the impact of tumor size variations.
5.4 Framework generalization
5.4.1 Effect of classifier selection
Table 3 compares different classifiers applied to the same set of encoder features. While linear probing and multilayer perceptron (MLP) classifiers, both implemented with default settings from scikit-learn, provide reasonable separability on the far-OOD cohorts, their performance degrades notably on the near-OOD cohorts (AUROC $\approx 92--93\%$, FPR95 $\approx 28--30\%$). In contrast, random forests achieve the best performance across all cohorts, highlighting the advantage of tree-based non-linear classifiers for capturing complex decision boundaries, particularly when distribution shifts are subtle (Figure 4(a)). The perfect performance on KiTS and PancreasCT (AUROC = 100%, FPR95 $\approx$ 0%) reflects the very easy nature of far-OOD detection when anatomical regions differ substantially from the training distribution.
5.4.2 Generalization to unseen OOD categories
Table 4: Generalization to unseen OOD categories across alternative training strategies on the random forests. Point estimates and 95% bootstrap confidence intervals over 100 matched-seed runs are reported.
Strategy AUROC $\uparrow$ FPR95 % $\downarrow$
RSNA PE MIDRC C19 KiTS PancreasCT RSNA PE MIDRC C19 KiTS PancreasCT
Dataset-specific
95.80
(92.60–98.50)
93.30
(89.90–96.00)
100.00
(99.90–100.00)
100.00
(100.00–100.00)
15.20
(7.10–23.50)
25.50
(16.80–36.20)
0.10
(0.00–1.00)
0.00
(0.00–0.00)
Ensemble
93.70
(91.70–96.50)
92.80
(89.80–95.60)
100.00
(99.80–100.00)
100.00
(100.00–100.00)
19.90
(11.20–29.10)
32.70
(20.40–54.10)
0.00
(0.00–0.00)
0.00
(0.00–0.00)
Unified
93.00
(85.50–97.50)
91.40
(85.40–94.90)
99.80
(99.00–100.00)
99.90
(99.30–100.00)
24.10
(11.20–42.90)
37.30
(20.90–69.40)
0.70
(0.00–5.60)
0.50
(0.00–4.10)
LODO
91.20
(85.50–95.30)
91.00
(84.40–94.80)
99.80
(99.30–100.00)
100.00
(99.70–100.00)
25.70
(17.30–36.70)
43.20
(23.50–54.10)
0.70
(0.00–4.10)
0.20
(0.00–2.00)
LODO+
92.10
(87.50–95.30)
91.80
(84.40–94.80)
99.80
(99.30–100.00)
100.00
(99.70–100.00)
21.70
(15.30–28.70)
34.20
(25.50–48.10)
0.70
(0.00–4.10)
0.20
(0.00–2.00)
We evaluated RF-Deep’s ability to generalize to previously unseen OOD categories using four additional training strategies (Table 4): (i) dataset-specific (DS; default), (ii) ensemble, (iii) unified, (iv) leave-one-dataset-out (LODO), and LODO+. DS trained a separate RF classifier for each OOD dataset, representing an upper-bound setting where the OOD type was known during development and evaluation. Ensemble trained the same four dataset-specific classifiers but averaged their predictions at test time, assuming no knowledge of the test domain during testing. Unified replaced the per-dataset training with a single RF classifier trained on ID versus all OOD datasets combined. LODO followed the standard OOD paradigm by training on ID plus three OOD datasets and evaluating on the held-out fourth, directly measuring generalization to novel OODs. Finally, LODO+ extended LODO by incorporating sub-sampled, non-tumor-centric ROIs to increase anatomical diversity during training.
Performance patterns on near-OOD datasets (RSNA PE and MIDRC C19) revealed important insights on within-anatomy generalization. The LODO strategy showed notable degradation compared to the dataset-specific baseline (RSNA PE: AUROC 91.20 vs 95.80, FPR95 25.70% vs 15.20%; MIDRC C19: AUROC 91.00 vs 93.30, FPR95 43.20% vs. 25.50%). This indicates that training on a single thoracic pathology provides insufficient diversity to generalize to pathologies occurring in the same anatomy. However, LODO+ improved performance over standard LODO (RSNA PE: FPR95 21.70% vs. 25.70%; MIDRC C19: FPR95 34.20% vs. 43.20%), demonstrating that increasing anatomical sampling diversity during training enhanced generalization to unseen near-OOD categories, which is critical for robust OOD detection.
The unified approach, which trained on all thoracic pathologies simultaneously, achieved intermediate performance between LODO and dataset-specific settings (RSNA PE: AUROC 93.00, MIDRC C19: AUROC 91.40), confirming that exposure to multiple within-anatomy pathologies during training improves cross-generalization within the same anatomical region. The ensemble strategy further improved near-OOD performance, approaching dataset-specific performance while maintaining the ability to detect novel OOD categories without prior knowledge of the test domain. On far-OOD datasets (KiTS and PancreasCT), all strategies achieved near-perfect performance (AUROC $\geq$ 99.80, FPR95 $\leq$ 0.70%), with minimal degradation from the dataset-specific baseline.
5.5 Design choices and sensitivity analysis
(a) Number of tumor-anchored ROIs
(b) Crop size of ROIs
(c) Number of training examples
(d) Individual stages of the Swin Transformer
Figure 9: Ablation studies showing the sensitivity of RF-Deep for OOD detection across four datasets. Point estimates and 95% bootstrap confidence intervals over 100 runs with matched seeds are displayed.
To understand the design choices of our proposed OOD detector, we conducted a series of controlled ablations on the number of ROIs, their crop sizes, the extent of labeled (supervised) data, and the encoder stages used for feature extraction. Figure 9(a) shows that OOD detection benefits from richer local context around predicted tumor regions, but only up to a point. With $n=1$, performance is unstable, reflected in wider confidence intervals and elevated FPR95, suggesting that a single tumor-anchored crop may miss informative variation in tumor surroundings. Increasing the number of ROIs from one to four improved performance on near-OOD cases, indicating that additional samples may better disambiguate similarly appearing tumor (ID) from non-cancerous structures. Beyond four ROIs, gains plateau and in some cases regress (e.g., MIDRC C19), implying diminishing returns and potential noise injection from redundant or low-value ROIs. Far-OOD datasets did not exhibit performance gains with increasing number of ROIs. Therefore, we used $n=4$ to capture meaningful contextual diversity while maintaining computational efficiency.
Figure 9(b) analyzes the effect of tumor-anchored 3D crop size on OOD detection performance. Performance remained relatively stable on the far-OOD cohorts across crop sizes, with slight preference for $128\times 128\times 128$. On the near-OOD cohorts, both smaller (<128<128) and larger (>128>128) ROIs led to performance degradation, suggesting that the optimal contextual window balances tumor-centered information with surrounding anatomical context.
Figure 9(c) shows that RF-Deep’s OOD detection performance improved with the number of labeled examples, using equal numbers of ID and OOD scans (5/5, 10/10, up to 20/20, corresponding to less than 30% of available data), after which performance plateaued. With only 20 ID and 20 OOD labeled scans, RF-Deep achieved AUROC >> 90 and FPR95 << 30% on average across all OOD datasets. Performance gains were more noticeable on the near-OOD cohorts than the far-OOD cohorts, indicating that detection on subtle shifts in the same anatomical region is trickier and requires more examples than larger shifts across different anatomical regions. Variance decreases consistently as more samples are added, with the strongest variance reduction observed on near-OOD cohorts.
Figure 9(d) shows individual Swin Transformer stage contributions. Far-OOD cohorts were relatively insensitive to individual stage selection (achieving AUROC >> 98 even when using features from a single stage), whereas the near-OOD counterparts showed strong sensitivity to early stages, with Stage 1 and Stage 2 features being particularly important (AUROC $\approx$ 88–90 individually). Concatenating features across all stages yielded the best performance, confirming that hierarchical representations are essential for robust OOD detection. This finding aligns with the SHAP analysis (Section 5.3.2), which showed that RF-Deep adaptively leverages features from different abstraction levels depending on the OOD scenario.
6 Discussion
We introduced a lightweight approach for instance-level OOD detection for lung tumor segmentation involving near-OOD and far-OOD cases. Our approach, termed RF-Deep, leveraged the advantage of SSL pretrained transformer encoders to extract image aquisition robust features, while using RF’s feature selection strategy to extract representations that distinguished subtle differences in pathologies occurring in the same disease site as the task model. Our approach improved performance over multiple current OOD detection methods as well as handcrafted radiomics-based RF classifiers under all evaluated datasets. We also found that RF-Deep approach was agnostic to the choice of pretraining approach as well as model size, thus providing a plug-and-play approach for leveraging an existing task-specific DL method for OOD detection.
Analysis of why RF-Deep outperformed other methods showed improved separability of ID from OOD datasets using deep compared to handcrafted radiomic features. Importantly, despite using the same deep features, Mahalanobis distance-based approach was less accurate than RF-Deep, indicating the need for outlier exposure and feature selection performed with random forests to extract a robust non-linear decision boundary. RF-Deep improved over logits-based methods, indicating that over-reliance on the ID model to extract confidence boundaries may be unreliable in OOD scans. Although this issue could be addressed to an extent through confidence calibration, prior studies show limitations of such methods (Larrazabal2023MEEP; mehrtash2020confidence).
Our approach has a few limitations. First, RF-Deep performs OOD detection only in scans where a segmentation was generated. Reduction of missing segmentations on ID scans is an altogether different problem and beyond scope of this work. Second, our approach was applied to binary tumor segmentation tasks. Extension to multi-class segmentation, with richer datasets, is straightforward and left for future work. Finally, the evaluated OOD datasets captured some of the key failure modes, but covering all possible pathologies is impractical due to lack of such datasets. Methodologies including continual learning with aggressive data augmentation for incorporating newly discovered failure modes are the next area of development.
7 Conclusion
We introduced RF-Deep, a post-hoc out-of-distribution (OOD) detection framework for segmentation models that leverages hierarchical encoder features and random forest classifiers to identify concept drifts. Across near-OOD and far-OOD cohorts, RF-Deep consistently outperformed logit-based methods and radiomics feature-based classifiers. The method is practical as it seamlessly integrates into existing segmentation pipelines without any modification, works across different pretraining strategies and architectures, and provides interpretable predictions through feature attribution.
Acknowledgment
This research was partially supported by the NCI R01CA258821 and the Memorial Sloan Kettering Cancer Center Support Grant/Core Grant NCI P30CA008748.
Appendix A Pretraining strategies descriptions
SMIT (jiang2022self) consists of masked voxel token prediction, masked patch token distillation, and image token distillation. Masked voxel token prediction is a dense regression task of the intensities within masked patches and is optimized by minimizing the $\ell_{1}$ norm between the predicted and ground-truth patches, akin to SimMIM (xie2022simmim). Self-distillation (antti2017meanteacher) was required for the next two tasks and was performed using an exponentially moving average (EMA) teacher model with identical architecture as the student, similar to iBOT (zhou2021ibot). Masked patch token distillation aligns the voxel token representations learned by the student with those extracted by the teacher from unmasked sequences, optimized with a temperature-scaled cross-entropy loss. For image token distillation, the cross-entropy between global volume embeddings of the student and teacher was minimized. After pretraining, the teacher was discarded and the student encoder was retained for fine-tuning.
Similarly, Swin UNETR (tang2022self) employs three pretraining tasks. The primary pretext task is masked volume reconstruction, where a proportion of 3D input patches are randomly masked and the model is trained to regress their voxel intensities from visible context. Unlike SimMIM, which leverages positional encodings to aid masked token prediction, this reconstruction task provided no explicit positional information, forcing the model to learn semantically plausible inpainting from local context. 3D rotation prediction enforces invariance by classifying the degree of rotation applied along the $z$-axis. Finally, contrastive coding encouraged representation learning by maximizing cosine similarity between augmented views of the same volume while minimizing similarity for different volumes. Unlike SMIT, which requires an EMA teacher, Swin UNETR used only a student model. The pretrained encoder was directly retained for fine-tuning.
Appendix B Logit-based approaches formulae
The tumor segmentation problem can be broken down into a per-voxel binary classification problem, in which we mean-aggregate the scores of the positive class to summarize the scan. In-distribution (ID) scans tended to yield more confident outputs, whereas out-of-distribution (OOD) scans produced less confident or spurious predictions. Following the convention of treating OOD as the positive class (hendrycks17baseline; choi2023conservative), we negated each score so that higher values correspond to greater likelihood of being OOD. Let $f^{i}(x)$ denote the logit for class $i\in\{0,1\}$ at voxel $x$, and define the softmax score as
$S_{\text{softmax}}^{i}(x)\;=\;\frac{\exp\!\big(f^{i}(x)\big)}{\sum_{j=0}^{1}\exp\!\big(f^{j}(x)\big)}.$
We denote the set of voxels predicted as positive (tumor) by
$V_{+}\;=\;\bigl\{(h,w,d)\,\big|\,\hat{y}_{h,w,d}=1\bigr\},$
where $\hat{y}_{h,w,d}$ is the class prediction at voxel $(h,w,d)$. Each logit-based score aggregates over $V_{+}$:
  • 1.
    MaxSoftmax (hendrycks17baseline):
    $S\;=\;-\,\frac{1}{|V_{+}|}\sum_{(h,w,d)\in V_{+}}S_{\text{softmax}}^{1}\!\left(x_{h,w,d}\right).$
  • 2.
    MaxLogit (hendrycks2022logits):
    $S\;=\;-\,\frac{1}{|V_{+}|}\sum_{(h,w,d)\in V_{+}}f^{1}\!\left(x_{h,w,d}\right).$
  • 3.
    Energy (liu2020energy):
    $S\;=\;-\,\frac{1}{|V_{+}|}\sum_{(h,w,d)\in V_{+}}\log\!\Bigg(\sum_{i=0}^{1}\exp\!\big(f^{i}(x_{h,w,d})\big)\Bigg).$
Appendix C Additional results
This section provides additional quantitative and qualitative results that support the findings presented in the paper.
  • 1.
    Figure 10 presents the disaggregated analysis of segmentation performance across imaging variations, discussed briefly in Section 5.1 (Figure 3).
  • 2.
    Table 5 lists the mapping from feature IDs used across Sections 5.2 and 5.3 to standardized IBSI feature names (retaining only those post recursive feature elimination).
  • 3.
    Figure 11 supports the discussion in Section 5.2 on MD-Deep, demonstrating that the standard formulation outperforms all variants.
  • 4.
    Table 6 reports the complete OOD detection performance of RF-Deep across all six evaluated backbones, briefly discussed in Section 5.2.
  • 5.
    Table 7 summarizes OOD performance for all methods (five additional pretraining strategies on six methods), reinforcing the findings in Section 5.2.
  • 6.
    Figure 12 shows t-SNE embeddings for all alternative pretraining strategies, complementing Figure 4(a) and discussions in Sections 5.2 and 5.3.1.
Figure 10: Segmentation performance across backbones under imaging variations (scanner type, contrast, reconstruction kernel) on the ID dataset. DSC and HD95 are reported.
Table 5: Radiomics features definitions.
ID Feature ID Feature
General features
feat_5feat\_5 maxIntensityInterpReseg feat_6feat\_6 minIntensityInterpReseg
Shape features
shape_7shape\_7 max2dDiameterSagittalPlane shape_11shape\_11 volume
shape_12shape\_12 filledVolume shape_18shape\_18 surfToVolRatio
First-order features
first_1first\_1 min first_2first\_2 max
first_4first\_4 range first_10first\_10 entropy
first_12first\_12 energy first_13first\_13 totalEnergy
Gray-Level Co-occurrence Matrix (GLCM) features
glcm_8glcm\_8 jointEntropy_3D_StdDev glcm_18glcm\_18 jointAvg_3D_StdDev
glcm_36glcm\_36 sumEntropy_3D_avg glcm_37glcm\_37 sumEntropy_3D_Median
glcm_39glcm\_39 sumEntropy_3D_Min glcm_43glcm\_43 contrast_3D_StdDev
glcm_44glcm\_44 contrast_3D_Min glcm_47glcm\_47 invDiffMom_3D_Median
glcm_48glcm\_48 invDiffMom_3D_StdDev glcm_58glcm\_58 invDiff_3D_StdDev
glcm_59glcm\_59 invDiff_3D_Min glcm_60glcm\_60 invDiff_3D_Max
glcm_61glcm\_61 invDiffNorm_3D_avg glcm_63glcm\_63 invDiffNorm_3D_StdDev
glcm_68glcm\_68 invVar_3D_StdDev glcm_76glcm\_76 diffEntropy_3D_avg
glcm_77glcm\_77 diffEntropy_3D_Median glcm_78glcm\_78 diffEntropy_3D_StdDev
glcm_79glcm\_79 diffEntropy_3D_Min glcm_80glcm\_80 diffEntropy_3D_Max
glcm_81glcm\_81 diffVar_3D_avg glcm_82glcm\_82 diffVar_3D_Median
glcm_84glcm\_84 diffVar_3D_Min glcm_89glcm\_89 diffAvg_3D_Min
glcm_92glcm\_92 corr_3D_Median glcm_93glcm\_93 corr_3D_StdDev
glcm_95glcm\_95 corr_3D_Max glcm_98glcm\_98 clustTendency_3D_StdDev
Gray-Level Run-Length Matrix (GLRLM) features
glrlm_28glrlm\_28 runLengthNonUniformityNorm_3D_StdDev glrlm_68glrlm\_68 grayLevelVariance_3D_StdDev
glrlm_73glrlm\_73 runLengthVariance_3D_StdDev glrlm_78glrlm\_78 runEntropy_3D_StdDev
Gray-Level Size Zone Matrix (GLSZM) features
glszm_3glszm\_3 grayLevelNonUniformity_3D
Gray-Tone Difference Matrix (GTDM) features
gtdm_5gtdm\_5 strength_3D
Table 6: RF-Deep OOD detection performance across different backbone encoders. AUROC ($\uparrow$) and FPR95 ($\downarrow$) with 95% bootstrap confidence intervals are reported over 100 matched-seed runs.
Backbone RSNA PE MIDRC C19 KiTS PancreasCT
AUROC FPR95 % AUROC FPR95 % AUROC FPR95 % AUROC FPR95 %
SimMIM
93.70
(90.40–96.10)
17.60
(10.20–28.10)
93.20
(90.10–96.10)
26.50
(18.30–38.20)
99.90
(99.80–100.00)
0.10
(0.00–1.00)
99.90
(99.80–100.00)
0.60
(0.00–2.00)
iBOT
93.60
(91.70–96.80)
22.60
(12.20–42.40)
92.70
(89.30–96.30)
24.60
(14.70–36.70)
99.80
(99.30–100.00)
1.10
(0.00–4.10)
99.80
(98.90–100.00)
1.10
(0.00–7.70)
SMIT Lite
94.10
(90.10–96.10)
20.40
(11.20–34.30)
90.40
(86.70–94.10)
30.50
(20.40–41.80)
99.60
(99.30–99.90)
1.80
(0.50–3.10)
99.30
(98.80–99.90)
1.80
(0.00–4.10)
Swin UNETR
89.60
(85.10–92.90)
34.10
(20.40–51.10)
88.90
(84.70–92.20)
37.50
(25.00–52.60)
99.70
(99.20–99.90)
2.40
(0.00–5.10)
99.60
(99.00–99.90)
2.90
(0.00–6.10)
Swin UNETR
90.40
(84.70–94.10)
41.50
(25.00–58.20)
88.60
(83.50–92.30)
35.80
(24.50–50.00)
99.40
(98.60–99.90)
2.60
(0.00–6.10)
98.70
(97.00–99.80)
7.10
(1.00–18.90)
SMIT
95.80
(92.60–98.50)
15.20
(7.10–23.50)
93.30
(89.90–96.00)
25.50
(16.80–36.20)
100.00
(99.90–100.00)
0.10
(0.00–1.00)
100.00
(100.00–100.00)
0.00
(0.00–0.00)
Figure 11: Ablation of Mahalanobis distance-based variants for OOD detection using AUROC over 100 matched-seed runs.
Table 7: RF-Deep OOD detection performance across different pretrained backbone encoders. AUROC ($\uparrow$) and FPR95 ($\downarrow$) with 95% bootstrap confidence intervals are reported over 100 matched-seed runs.
Model Method RSNA PE MIDRC C19 KiTS PancreasCT
AUROC FPR95 % AUROC FPR95 % AUROC FPR95 % AUROC FPR95 %
SimMIM MaxSoftmax 86.89 45.56 84.58 54.92 94.19 21.59 94.30 24.04
MaxLogit 89.39 44.04 87.55 45.65 95.99 15.32 96.28 14.82
Energy 89.70 44.64 87.75 45.24 96.15 14.92 96.39 13.84
RF-Radiomics 87.20 48.90 89.80 34.90 97.00 15.80 96.60 17.90
MD-Deep 88.60 41.32 86.85 51.72 97.66 3.40 97.84 8.30
RF-Deep 93.70 17.60 93.20 26.50 99.90 0.10 99.90 0.60
iBOT MaxSoftmax 83.56 48.25 80.97 52.10 90.45 32.22 88.98 36.23
MaxLogit 87.31 38.55 85.24 39.33 93.82 23.45 92.29 27.17
Energy 87.90 38.62 85.84 39.61 94.49 22.02 92.96 25.17
RF-Radiomics 89.30 50.90 90.50 29.70 96.80 14.70 96.80 15.20
MD-Deep 87.83 44.87 86.12 52.89 98.63 6.62 96.16 17.27
RF-Deep 93.60 22.60 92.70 24.60 99.80 1.10 99.80 1.10
SMIT Lite MaxSoftmax 77.96 56.05 74.44 60.96 80.77 48.42 87.35 41.45
MaxLogit 78.52 53.32 79.03 51.95 79.46 45.97 85.42 41.33
Energy 78.16 53.82 79.37 52.22 78.74 45.59 83.90 40.53
RF-Radiomics 88.90 38.90 90.70 31.10 97.10 13.80 97.40 12.80
MD-Deep 85.94 38.85 78.51 56.77 93.51 25.57 92.26 25.06
RF-Deep 94.10 20.40 90.40 30.50 99.60 1.80 99.30 1.80
Swin UNETR MaxSoftmax 80.30 52.19 79.65 62.01 84.40 46.38 88.97 46.98
MaxLogit 83.88 42.62 83.85 52.12 88.99 34.86 91.57 35.45
Energy 84.16 40.77 84.30 50.44 89.41 31.86 92.17 33.91
RF-Radiomics 87.70 37.40 89.20 33.40 98.10 9.80 98.10 13.70
MD-Deep 77.88 62.47 74.14 64.61 92.61 25.50 89.54 29.67
RF-Deep 89.60 34.10 88.90 37.50 99.70 2.40 99.60 2.90
Swin UNETR MaxSoftmax 75.92 53.25 77.83 55.91 78.89 53.17 93.12 23.96
MaxLogit 77.87 43.65 80.41 49.93 83.33 39.30 92.37 18.44
Energy 78.02 42.46 80.78 49.48 83.87 38.24 92.17 18.98
MD-Deep 74.03 68.07 70.50 71.02 84.56 49.44 79.45 63.34
RF-Radiomics 86.50 40.00 89.00 37.20 97.40 11.30 97.80 11.70
RF-Deep 90.40 41.50 88.60 35.80 99.40 2.60 98.70 7.10
(a) SimMIM
(b) iBOT
(c) SMIT Lite
(d) Swin UNETR
(e) Swin UNETR
Figure 12: t-SNE projected embeddings showing dataset-wise separability of (a) deep features and (b) radiomics features. Results from one representative split (of 100), combining all datasets in a single visualization, are shown for brevity. The shaded blue regions indicate the convex hull of the ID dataset.