### v2

# Tumor-anchored deep feature random forests for out-of-distribution detection in lung cancer segmentation

Abstract
Accurate segmentation of cancerous lesions from 3D computed tomography (CT) scans is essential for automated treatment planning and response assessment. However, even state-of-the-art models combining self-supervised learning (SSL) pretrained transformers with convolutional decoders are susceptible to out-of-distribution (OOD) inputs, generating confidently incorrect tumor segmentations, posing risks to safe clinical deployment. Existing logit-based methods suffer from task-specific model biases, while architectural enhancements to explicitly detect OOD increase parameters and computational costs. Hence, we introduce a lightweight, plug-and-play post-hoc random forests-based OOD detection framework called RF-Deep that leverages deep features with limited outlier exposure. RF-Deep enhances generalization to imaging variations by repurposing the hierarchical features from the pretrained-then-finetuned backbone, providing task-relevant OOD detection by extracting the features from multiple regions of interest anchored to the predicted tumor segmentations. We compared RF-Deep against existing OOD detection methods using 2,056 CT scans across near-OOD (pulmonary embolism, negative COVID-19) and far-OOD (kidney cancer, healthy pancreas) datasets. RF-Deep achieved AUROC >> 93.50 for the challenging near-OOD datasets and near-perfect detection (AUROC >> 99.00) for the far-OOD datasets, substantially outperforming logit-based and radiomics approaches. RF-Deep maintained consistent performance across networks of different depths and pretraining strategies, demonstrating its effectiveness as a lightweight, architecture-agnostic approach to enhance the reliability of tumor segmentation from CT volumes.
1 Introduction
Deep learning (DL) based medical image analysis is now applicable throughout the continuum starting from diagnosis, treatment planning, and response monitoring at follow-up to enhance care of patients with cancer. For instance, hybrid transformer-convolutional networks have been shown to provide highly accurate tumor and healthy tissue segmentations from radiographic modalities (Jiang et al., 2022; Willemink et al., 2022; Nguyen et al., 2023; Yan et al., 2023; Qayyum et al., 2023; Gu et al., 2025). However, these methods are typically developed on narrowly-scoped datasets, and accuracy on in-distribution (ID) data does not translate to out-of-distribution (OOD) scenarios including different imaging acquisitions (Roschewitz et al., 2023; Koch et al., 2024), concept drifts (Huang et al., 2022; Sahiner et al., 2023; Gomez et al., 2025), and label shifts (Nichyporuk et al., 2022; Lempart et al., 2023; Godau et al., 2025). Importantly, DL models produce confidently inaccurate predictions (Hendrycks et al., 2019; Quiñonero-Candela et al., 2022; Prabhu et al., 2022; Banerjee et al., 2023), preventing clinicians from reliably using model uncertainties for decision making (Otles et al., 2021) and risking clinical expert deskilling through over-reliance on AI (Budzyń et al., 2025).
Common deployment failures that motivate our work include the Picture Archiving and Communication System (PACS) routing errors directing non-thoracic scans to lung-specific analysis pipelines and rare thoracic pathologies (e.g., pulmonary embolism) mimicking tumor appearances. Segmentation accuracy metrics like Dice and Hausdorff distances do not reflect well-calibrated uncertainty estimates, even on ID datasets (Ren et al., 2024; Rangnekar et al., 2025). While uncertainty visualizations can enhance clinical delineations (Goddard et al., 2012; Biase et al., 2024), such measures are unreliable in OOD scenarios, as shown in Figure 1. Existing OOD detection methods are often too sensitive, flagging any deviation from training data (Roschewitz et al., 2023; Godau et al., 2025), limiting their use for practical purposes.
Our work overcomes the limitations of prior approaches by leveraging pretrained models to ensure sufficient generalization to common clinical imaging variations while enhancing robustness to far-OOD (disease in different anatomical sites) and near-OOD (different disease in same anatomic site) use cases. We propose a tumor-anchored deep feature extraction framework for post-hoc OOD detection. Unlike prior logit-based or distance methods that operate on entire images or require architectural modifications, we extract hierarchical encoder features from multiple regions of interest anchored to predicted tumor segmentations. This enables task-relevant detection that distinguishes subtle pathological differences within the same anatomical site while scaling to images with varying fields-of-view. Our contributions are:
Figure 1: Uncertainty maps for in-distribution (ID) and out-of-distribution (OOD) scans. (a) ID lung cancer scans (1–2) show concentrated boundary uncertainty; OOD scans (3–6: pulmonary embolism, COVID-19-negative, kidney cancer, healthy abdomen) exhibit concentrated, diffused, or misaligned patterns.
  • A post-hoc OOD detection approach using random forests with deep features (RF-Deep), trained with outlier exposure. Features are extracted from regions of interest (ROIs) anchored to the task-specific model’s predictions, enabling scan-level and task-relevant OOD filtering. Our approach reduced false positive detections, achieving AUROC >> 90 on challenging near-OOD datasets and near-perfect detection on the far-OOD datasets.
  • Generalizability to common imaging variations is ensured by using features extracted from the hierarchical encoder layers of a segmentation model subject to self-supervised learning (SSL) pretraining and task-specific fine-tuning. Robustness of task-specific model performance on commonly encountered imaging variations is also reported.
  • Evaluation on 2,056 3D CT scans from five public datasets, with analysis of SSL pretraining methods, model depths, and classifier designs on OOD detection performance.
  • Analysis of why RF-Deep outperforms radiomics-based and logit-based approaches by performing unsupervised clustering, SHAP-based feature importance, and spatial uncertainty analysis.
2 Related Works
Approaches for OOD detection fall into two classes: learner-based methods that extract distribution-robust features through training strategies, and filter-based methods that explicitly flag OOD scans at test time.
2.1 Learner-based methods
Learner-based methods extract robust feature representations from available training data so models generalize across domains without performance degradation. Domain-invariant features can be obtained via (a) specialized architectures (Saha et al., 2021; Simeth et al., 2023), (b) data augmentation in image and feature space (Jung et al., 2019; Chen et al., 2022; Vaish et al., 2025), (c) regularized training with multi-task objectives (Panfilov et al., 2019; Wen et al., 2024), and (d) SSL pretraining on large unlabeled examples (Hatamizadeh et al., 2022; Haghighi et al., 2022; Jiang et al., 2024; Du et al., 2024).
For tumor segmentation, architectures combining attention mechanisms and residual connections have achieved robust generalization across MR acquisitions (Saha et al., 2021; Simeth et al., 2023). Gu et al. (2021) trained separate domain-specific basis functions combined via group convolutions for prostate gland segmentation. Alternatively, data augmentation (Safdari et al., 2025) and regularization (Jung et al., 2019; Vaish et al., 2025) have been applied to standard architectures like U-Net (Chen et al., 2022; Wen et al., 2024) to further enhance robustness to distribution shifts (Boone et al., 2023).
SSL pretraining takes a different approach, exploiting unlabeled data through extensive data augmentations (Fedorov et al., 2021) or exposure to diverse imaging variations (Hatamizadeh et al., 2022; Jiang et al., 2024; Gomez et al., 2025). Azizi et al. (2023) proposed successive refinement, converting OOD data to ID through iterative adaptation, although unlike SSL, this requires labeled data for each new domain.
2.2 Filter-based OOD methods
Filter-based methods reject OOD scans at test time, ensuring models operate only on ID-like data where predictions are reliable (Berger et al., 2021). These methods implement either separate OOD detectors or task-integrated detection, and have been applied to handle imaging acquisition differences (Tardy et al., 2019; Gao and Wu, 2020; Nandy et al., 2021). Most filter-based methods focused on 2D classification, with approaches ranging from Mahalanobis distances (Tardy et al., 2019) to dedicated OOD networks (Nandy et al., 2021) to instance-level retraining that measured response variations on ID test sets (Gao and Wu, 2020). A different approach involves using a single model to classify tasks on inlier data, while exposing the same model to outlier examples to detect OOD during inference (Roy et al., 2022; Araujo et al., 2023). This approach is practical when the OOD types are known from domain knowledge (e.g., optical tomography versus retinal fundus images) (Araujo et al., 2023), or when outliers such as infrequent disease conditions are identifiable from the training data (Roy et al., 2022).
For segmentation, Mahalanobis distance on feature embeddings has been applied to COVID-19 lung lesion detection (González et al., 2022), but the high dimensionality of feature embeddings limits accuracy, particularly for distinguishing pathologies within the same disease site (Berger et al., 2021; Roy et al., 2022). Dimensionality reduction via PCA (Woodland et al., 2024) or VQ-GAN (Pinaya et al., 2022; Graham et al., 2023) has been used to reduce the latent space and sharpen the separation of ID from OOD feature distributions, but introduces inductive bias (Nalisnick et al., 2019) and requires additional training. Logit-based scores (Hendrycks and Gimpel, 2017; Hendrycks et al., 2022) and temperature scaling (Karimi and Gholipour, 2022) offer computational simplicity but inherit task-model biases (Mehrtash et al., 2020). Architectural modifications (Yuan et al., 2023; Larrazabal et al., 2023) and generative diffusion methods (Nguyen et al., 2026) have improved OOD detection but increase network parameters and training complexity. Radiomic feature distances have shown utility for dataset-level OOD detection (Vasiliuk et al., 2023; Konz et al., 2024) but struggle with scan-level detections, as they cannot simultaneously capture subtle data variations while providing robust performance to common imaging variations.
Figure 2: RF-Deep workflow for scan-level OOD detection. Panels (a–c) depict feature extraction from the frozen Refer to caption segmentation model and random forest training in an outlier exposure manner and panel (d) shows scan-level ID/OOD inference using the trained detector.
3 Framework overview and OOD task definition
To address these limitations, our goal was to develop a highly accurate scan-level OOD detection approach, which does not require large numbers of outlier samples, with generalizable applicability to images with different acquisitions and fields of view. Let $\mathcal{D}_{\text{in}}$ denote the distribution of lung cancer CT scans, closely matching the training dataset used to create the lung tumor segmentation model, and let $\mathcal{D}_{\text{out}}$ represent the scans that differ in pathology and anatomic sites. Given a new scan $x$, the objective is to determine whether it belongs to $\mathcal{D}_{\text{in}}$ or $\mathcal{D}_{\text{out}}$ at the scan level using a scoring mechanism provided by a random forest (RF) classifier.
Our approach combines the strengths of learner-based and filter-based OOD detection methods. Specifically, we leverage the backbone of a segmentation model pretrained via self-supervised learning to extract imaging features that are largely domain-agnostic, and pair these features with a lightweight random forest classifier to reject conceptually dissimilar scans. The RF-based OOD classifier incorporates known outliers during training, an outlier exposure strategy shown to be effective in prior work (Hendrycks et al., 2019; Thulasidasan et al., 2021; Guha Roy et al., 2022). We adopt outlier exposure for three key reasons. First, in clinical deployment, common OOD failure modes (e.g., incorrect body region or scanner-related differences) are often predictable based on domain knowledge. Second, purely unsupervised methods (such as Mahalanobis distance–based approaches) frequently struggle with near-OOD scans (Table 1), where pathologies share anatomical context with the in-distribution data. Third, our method requires only a small number of outlier examples (as few as 20 scans) to achieve strong performance (Section 5.5), making outlier curation feasible in practice.
Importantly, we evaluate generalization to unseen OOD categories via leave-one-dataset-out experiments (Section 5.4.2), demonstrating that RF-Deep extends beyond the specific outliers encountered during training. Scan-level OOD prediction for segmentation involves thousands of voxel-level predictions, which, while offering rich spatial context, also introduces memory constraints in processing and aggregating information from their large 3D volumes. Our approach addresses this issue by focusing the OOD detection in regions of interest (ROI) anchored to the predicted tumor segmentation, and then averaging the metric scores from those regions, which also allowed the approach to scale to images with varying fields of view. Using regions anchored near the predicted tumors also allows the model to focus on differentiating features in relevant parts of the image, as opposed to distinguishing entire images, enabling distinction of pathologies occurring in the same disease site.
A key design consideration is RF-Deep’s dependence on the segmentation model’s predictions. We observed that on OOD scans, the segmentation model typically produced confidently incorrect tumor-like regions, rather than failing to produce any output. These predictions served as anchors for ROI extraction, and the extracted features from these regions are precisely what RF-Deep uses to identify the scan as OOD. In rare cases where the model produces no foreground predictions, the absence itself can serve as an implicit OOD signal, as ID scans with confirmed tumors would be expected to yield predictions. Across our evaluation datasets, the segmentation model produced at least one connected component on almost 90% of OOD scans, confirming that the tumor-anchored approach remains applicable even under distribution shift.
3.1 Segmentation model
The segmentation framework employs a hybrid transformer-convolution architecture to leverage strengths of global contextual modeling with the transformer backbone as encoder and local spatial precision of the convolutional decoder. The backbone is a hierarchical Swin Transformer (Liu et al., 2021) with depth configuration 221222-2-12-2 across four successive stages, with a patch size of $2\times 2\times 2$ voxels and a window size of $4\times 4\times 4$ voxels. Its hierarchical design progressively aggregates features at multiple spatial scales to capture fine-grained anatomical details. Windowed self-attention within each stage reduces complexity while preserving details and global context, critical for processing large volumetric images of size $128\times 128\times 128$ voxels. The backbone was pretrained with the self-distilled masked image transformer (SMIT) framework to enhance robustness to imaging and patient variations (Jiang et al., 2022). SMIT combines masked image prediction (He et al., 2021; Xie et al., 2022) with self-distillation (Zhou et al., 2022); described in Appendix A. The decoder, a 3D U-Net-based (Ronneberger et al., 2015) convolutional structure, was randomly initialized and fine-tuned on the ID task-specific dataset. Following training, the entire model is frozen ([Uncaptioned image]) for subsequent segmentation and feature-based OOD detection.
3.2 Scan-level OOD detection approach
Our OOD detection pipeline consists of the following four steps (Figure 2):
  • Regions of interest extraction: The task-specific segmentation model (Subsection 3.1) was used to generate tumor segmentations from the entire scan. Detected tumor regions were used to extract multiple 3D ROIs via random spatial offsets, ensuring the tumor appears within each ROI without precise centering (Figure 2a).
  • Feature extraction. We repurposed the fine-tuned backbone encoder from the segmentation model to extract multi-scale features from the five encoder stages: the patch embedding layer (48 channels) and the four Swin Transformer stages (96, 192, 384, and 768 channels, respectively). All features were aggregated into a single scan-level descriptor through global average pooling (GAP), producing a single feature vector per scan (Figure 2b).
  • Training the OOD detector. Feature vectors were extracted from ID lung cancer datasets as well as a representative set of OOD scans from tumor-anchored regions as described above and used to train the RF classifier. This approach enabled our detector to learn subtle decision boundaries that distinguish shifts even between different diseases affecting the same anatomic region (Figure 2c).
  • Online inference. Given a new test instance, the aforementioned pipeline extracts potential tumor regions and tumor-anchored ROIs, followed by the extraction of GAP features, which the RF classifier uses to generate scan-level OOD classification (Figure 2d).
4 Experimental setup
4.1 Implementation details
All models were implemented in PyTorch (Paszke et al., 2019) and MONAI (Cardoso et al., 2022). Our method operates on full 3D CT volumes at full resolution, resampled to a uniform isotropic voxel spacing of $1~\text{mm}^{3}$. All scans were clipped to the lung window (Hounsfield units (HU): [400,400][-400,400]) and then intensity-normalized. During inference, the segmentation model processes the entire volume using sliding window inference ($128\times 128\times 128$ voxel patches with 50% Gaussian overlap) to generate tumor segmentations. For OOD detection, every detected tumor seeds multiple tumor-anchored 3D ROIs ($128\times 128\times 128$ voxels each) around the predicted tumor areas; the OOD scores from these regions are combined to generate a scan-level prediction.
Segmentation. Models were fine-tuned using cross-entropy and Dice loss with a batch size of 16 distributed across four NVIDIA GPUs, with an input size of 128 $\times$ 128 $\times$ 128 voxels. We used AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of $2\times 10^{-4}$, linear warm-up (50 epochs), and cosine annealing over 1,000 epochs. Data augmentations included random flips, rotations, affine transformations, and illumination adjustments, consistent with Tang et al. (2022); Jiang et al. (2022).
OOD Detection. Random forest classifiers employed 1,000 trees with a maximum depth of 20 (using default scikit-learn hyperparameters). We extracted deep feature representations from $n=4$ tumor-anchored 3D ROIs per scan for all ID and OOD datasets. All ROIs were used individually during training to increase training samples, and probabilities from all ROIs were averaged at the scan level as an ensemble for inference. For MaxSoftmax, MaxLogit, and energy, scan-level scores were obtained by mean-aggregating tumor (positive class) predictions across all voxels using the corresponding metric.
4.2 Datasets
Backbone pretraining. The backbone encoder was pretrained using the self-distilled masked image transformer (SMIT) framework on 10,432 3D CT scans, encompassing public and institutional datasets, following Jiang et al. (2022); Tang et al. (2022); Jiang et al. (2024).
Comparison to other pretraining strategies and architectures. Additionally, we pretrained using SimMIM (Xie et al., 2022), iBOT (Zhou et al., 2022), and Swin UNETR (Tang et al., 2022) to assess the impact of pretraining strategy on both accuracy and OOD robustness. To analyze the effect of model depth, we pretrained a lite backbone configuration (22222-2-2-2) using SMIT. Finally, we evaluated the publicly available Swin UNETR pretrained checkpoint (marked as ) to assess the impact of pretraining data.
We do not consider interactive foundation models for segmentation, such as SAM (Kirillov et al., 2023) or its medical variants (e.g., MedSAM (Ma et al., 2024) and VISTA3D He et al. (2025)), as these methods are primarily designed for prompt-guided segmentation and require user-provided inputs (points or bounding boxes), which precludes fully automated deployment. Additionally, we exclude self-supervised 2D models such as DINOv2 (Oquab et al., 2023), which are pretrained on natural images and operate on individual slices, thereby losing critical inter-slice spatial context necessary for distinguishing 3D pathological structures in CT volumes (that is, relying on 2.5D processing).
Fine-tuning. We combined the pretrained backbone as the encoder with a randomly initialized U-Net decoder and fine-tuned using $N=317$ 3D scans from the publicly available non-small cell lung cancer (NSCLC) Radiomics dataset (Aerts et al., 2015) with radiologist-provided tumor delineations. The original dataset contains 422 scans; 317 were selected based on radiologist assessment and verification of tumor boundary quality.
Testing. We used the NSCLC Radiogenomics dataset (Bakr et al., 2017) as the ID test set. From the original 211 scans, $N=140$ scans from various scanners (GE, Siemens) were selected based on radiologist assessment and verification of tumor boundaries and image quality. The ID test set is completely separate from the training set, acquired from different patients and lung cancer stages, providing external validation. OOD evaluation used 1,916 3D CT scans from four public datasets: pulmonary embolism (RSNA PE, $N=1,225$ selected from 3,114 scans, preserving the marginal distribution of labels in the full PE-positive dataset) (Colak et al., 2021), COVID-19 negative patients (MIDRC C19, $N=120$, all used) (Tsai et al., 2021), kidney cancer (KiTS, $N=489$, all used) (Heller et al., 2023), and healthy scans of the pancreas (PancreasCT, $N=82$, all used) (Roth et al., 2015). The near-OOD datasets (RSNA PE, MIDRC C19) consist of chest CTs involving non-cancerous pathologies such as pulmonary embolism and inflammation, depicting appearances similar to lung lesions. The far-OOD datasets (KiTS, PancreasCT) include scans with larger field-of-view than training with healthy organs and non-lung cancers. Patient-level splits ensured zero overlap between training and evaluation; OOD counts were held constant across runs.
4.3 Evaluation metrics
Segmentation accuracy on ID datasets was measured using the Dice similarity coefficient (DSC) and the 95th percentile Hausdorff distance (HD95). OOD detection accuracy was computed using the area under the receiver operating characteristic curve (AUROC) and false positive rate at 95% true positive rate (FPR at 95%TPR, or FPR95). AUROC quantifies the separability between ID and OOD scans (Davis and Goadrich, 2006), with higher values indicating better discrimination. FPR95 measures the proportion of OOD scans misclassified as ID when sensitivity is fixed at 95% (Liang et al., 2018), where lower values are better.
To ensure meaningful AUROC computation, the number of OOD scans was balanced with ID test scans ($N=140$) by subsampling OOD examples without replacement from larger datasets (RSNA PE and KiTS) and averaging the computed metrics. Splits ensured zero patient overlap with consistent OOD scan counts. This evaluation approach avoids biases from imbalanced evaluation sets while ensuring stable ID-OOD ratios, following best practices for robust OOD evaluation (Szyc et al., 2023; Humblot-Renaux et al., 2023).
For segmentation performance, we reported the mean and standard deviation across test scans. For OOD detection performance, we reported the point estimates with 95% bootstrap confidence intervals computed over 100 runs with matched random seeds to reflect sampling variability across different train-validation splits.
4.4 Comparative OOD detection methods
Baseline comparisons included commonly used OOD detection methods, namely, MaxSoftmax (Hendrycks and Gimpel, 2017), MaxLogit (Hendrycks et al., 2022), and energy (Liu et al., 2020), performed by computing scan-level scores by mean-aggregating the tumor (positive class) predictions across all voxels (see Appendix B).
In addition, a RF classifier was created using handcrafted radiomics features ($N=293$) called RF-Radiomics, computed within the segmented tumor regions following the Image Biomarker Standardization Initiative (IBSI) guidelines (Zwanenburg et al., 2020) with open-source software PyCERR (Apte et al., 2018). Critically, RF-Radiomics uses identical outlier exposure as RF-Deep (same ID and OOD training splits), isolating the effect of feature representation: any performance difference between RF-Deep and RF-Radiomics is attributable to deep features versus handcrafted features, not to differential access to OOD supervision. Details of individual radiomic features are in Table 5. The high correlation among radiomic features was addressed by employing recursive feature elimination to select a more compact and informative set of radiomic features. Of note, recursive feature elimination was not used for training the RF-Deep classifier.
Finally, the Mahalanobis distance-based OOD detector (MD-Deep) (Lee et al., 2018) was computed with respect to deep features of the ID data, modeled as a single Gaussian in feature space using Ledoit-Wolf covariance estimation. Hence, MD-Deep computes OOD scores as the distance of each scan with respect to the fitted ID distribution. MD-Deep uses identical backbone features as RF-Deep but operates without outlier exposure, isolating the contribution of OE and the RF classifier’s learned decision boundaries. Comparing RF-Deep to MD-Deep thus reveals the benefit of outlier exposure and non-linear classification over purely unsupervised density estimation on the same feature space. We excluded ODIN (Liang et al., 2018) because it relies on temperature scaling and input perturbations calibrated on fine-tuning data, parameters that are often unavailable when deploying models to new out-of-distribution clinical datasets.
(a) Segmentation metrics
(b) Imaging variations
(c) Segmentation examples
Figure 3: Pretraining strategies performance and robustness evaluation on ID test set. (a) Segmentation metrics (DSC and HD95). (b) Performance across imaging variations. (c) Representative segmentations; R denotes reconstruction kernel.
5 Results
5.1 Segmentation performance on ID dataset
Segmentation performance for the various analyzed pretraining strategies is shown in Figure 3(a) for the ID testing dataset. All pretrained methods were similarly accurate, with iBOT producing the lowest accuracy, followed by the publicly available Swin UNETR checkpoint.
Next, we performed a disaggregated analysis of model performance under common imaging variations including scanner manufacturer (GE versus non-GE), contrast versus non-contrast enhanced CT, and reconstruction kernels. Reconstruction kernels were categorized as Recon 1 (GE ‘standard’ and ‘bone’, Siemens with << B40), Recon 2 (GE ‘Bone Plus’, Siemens with $\geq$ B40 and << B50), and Recon 3 (GE ‘Lung’, Siemens $\geq$ B50). Model fine-tuning was performed on the public dataset acquired on Siemens scanners with contrast and Recon 1 kernel. From Figure 3(b), all models were similarly accurate across the analyzed variations, with larger gaps in accuracy noted for Swin UNETR (Recon 1, non-contrast CT) and iBOT (Recon 3, non-contrast). All other strategies showed relatively small degradation in accuracy due to the analyzed variations indicating robust performance through any of the pretrained methods. SMIT and SMIT Lite strategies were similarly accurate, indicating that a smaller model is sufficient for robust tumor segmentation. On the other hand, choice of pretraining data impacted accuracy, as seen by the performance difference between the Swin UNETR trained with the 10K dataset and the public checkpoint Swin UNETR. Detailed accuracies using DSC and HD95 metrics for the analyzed methods are shown in Figure 9.
Overall, the SMIT-pretrained model was the most accurate, followed by the Swin UNETR method, and was thus selected for benchmarking the OOD detection method using RF-Deep throughout this paper.
5.2 Out-of-distribution detection performance
RF-Deep is the most accurate method for both near-OOD and far-OOD datasets (Table 1). On the near-OOD datasets (RSNA PE and MIDRC C19), RF-Deep achieved AUROCs of 95.80 and 93.30, outperforming the next best method (MaxLogit) by approximately 7.0 and 4.0 percentage points, respectively. Similarly, RF-Deep reduced FPR95 by 18–25 absolute points compared to the next closest competitor. These improvements remain consistent within the 95% confidence intervals, with non-overlapping intervals confirming statistically significant superiority over all baseline methods, highlighting the stability of RF-Deep across bootstrap resampling.
Table 1: OOD detection performance comparing various methods on the segmentation model fine-tuned using the SMIT-pretrained backbone. AUROC ($\uparrow$) and FPR95 ($\downarrow$) are reported with 95% bootstrap confidence intervals over 100 matched-seed runs.
Method RSNA PE MIDRC C19 KiTS PancreasCT
AUROC FPR95 % AUROC FPR95 % AUROC FPR95 % AUROC FPR95 %
MaxSoftmax
88.61
(84.62–92.36)
38.37
(31.78–49.24)
86.57
(81.79–90.14)
49.27
(39.53–59.78)
95.67
(93.41–98.06)
23.26
(11.47–34.19)
94.61
(92.61–97.49)
34.27
(24.80–49.30)
MaxLogit
88.77
(85.09–92.43)
40.01
(24.80–52.52)
89.31
(85.05–93.06)
41.59
(29.84–53.96)
95.89
(93.58–98.44)
18.05
(10.41–27.37)
93.53
(90.17–96.59)
30.87
(16.17–56.87)
Energy
88.59
(84.85–92.24)
39.86
(24.80–52.55)
89.30
(84.98–93.04)
43.04
(29.84–55.43)
95.80
(93.41–98.44)
17.47
(10.41–26.65)
93.52
(90.10–96.53)
29.96
(15.83–56.87)
RF-Radiomics
88.70
(81.50–93.00)
41.10
(25.20–57.80)
91.30
(87.10–94.40)
30.00
(17.40–45.60)
96.20
(93.00–98.30)
11.80
(5.80–21.90)
96.10
(89.10–99.20)
17.20
(4.30–41.90)
MD-Deep
87.30
(83.90–91.10)
45.10
(30.60–60.30)
84.50
(79.40–89.00)
58.50
(38.20–77.60)
99.20
(97.90–99.80)
3.10
(0.00–9.20)
99.20
(98.20–100.00)
2.80
(0.00–7.70)
RF-Deep
95.80
(92.60–98.50)
15.20
(7.10–23.50)
93.30
(89.90–96.00)
25.50
(16.80–36.20)
100.00
(99.90–100.00)
0.10
(0.00–1.00)
100.00
(100.00–100.00)
0.00
(0.00–0.00)
RF-Deep also achieved near-perfect separability on the far-OOD datasets using KiTS (AUROC 100, FPR95 0.10%) for kidney cancers and PancreasCT (AUROC 100, FPR95 0.00%) containing healthy abdominal scans, effectively eliminating false alarms for anatomically distant pathologies. Although it was less accurate on the near-OOD datasets compared to its accuracy on the far-OOD datasets, the gap in AUROC (RSNA PE: ++3.5%, MIDRC C19: ++6.7%) was the smallest compared to all other methods, with the next smallest gap observed for RF-Radiomics (RSNA PE: ++7.5%, MIDRC C19: ++4.9%).
RF-Radiomics, while competitive on the far-OOD datasets (AUROC of 96.20 on KiTS and 96.10 on PancreasCT with relatively low FPR95), produced moderate error rates on the near-OOD datasets (AUROC 88.70 and 91.30, FPR95 3041%), indicating that handcrafted features are less effective than deep features at capturing subtle differences in the appearance of pathologies (analyzed further in Sections 5.3.1 and 5.3.2).
MD-Deep, although highly accurate for detection on far-OOD datasets (AUROC 99.2 on both KiTS and PancreasCT) was less effective and similar to logit-based methods on near-OOD datasets (AUROC 87.3 on RSNA PE and 84.5 on MIDRC C19, with FPR95 around 45-59%). MD-Deep was also less accurate than RF-Deep on near-OOD datasets, despite both methods using identical encoder features, indicating that feature selection performed by the RF was essential to achieve better separation of ID from OOD scans. Additional evaluation to assess if feature reduction enhanced performance showed no improvement compared to the standard MD-Deep approach (Figure 10).
All logit-based methods showed a considerable drop in accuracy for the near-OOD datasets compared to their performance on far-OOD datasets, reflecting the challenge in detecting OOD from images depicting different pathologies but from the same disease site as the ID dataset. For instance, MaxLogit exhibited performance gaps of 7–8% between near-OOD and far-OOD datasets, more than twice the gap observed for RF-Deep (3.5–6.7%), suggesting that logit-based scores alone may be insufficient for challenging near-OOD detection. We analyzed the underlying causes of these performance differences in Section 5.3.3.
The controlled comparisons isolate specific contributions: (1) RF-Deep vs. RF-Radiomics (same classifier, same OE, different features) demonstrates that deep features outperform handcrafted radiomics, particularly on near-OOD where subtle pathological differences matter; (2) RF-Deep vs. MD-Deep (same features, different classifier, with vs. without OE) demonstrates that outlier exposure combined with RF’s non-linear decision boundaries outperforms unsupervised density estimation; (3) RF-Deep vs. logit-based methods (different paradigm entirely) demonstrates that encoder features are more discriminative than output-layer uncertainties for OOD detection.
5.2.1 Impact of model pretraining on OOD detection
Next, we analyzed whether the choice of model pretraining, model size (SMIT versus SMIT Lite), and pretraining data (Swin UNETR versus Swin UNETR) impacted the accuracy of RF-Deep. RF-Deep remained the most accurate method across all backbone architectures (Table 6), consistently achieving near-perfect accuracy on the far-OOD datasets. Table 7 further demonstrated that RF-Deep’s superiority over logit-based and radiomics methods was consistent across all tested backbones, with RF-Deep outperforming the next best method by large margins on the near-OOD datasets. Accuracy differences between SMIT and SMIT Lite models on near-OOD datasets were modest, with SMIT Lite showing slightly better performance on RSNA PE but marginally lower on MIDRC C19.
Similarly, RF-Deep performance for Swin UNETR and Swin UNETR backbones was comparable (RSNA PE AUROC: 90.4 vs 89.6; MIDRC C19 AUROC: 88.6 vs 88.9; RSNA PE FPR95: 41.5% vs 34.1%; MIDRC C19 FPR95: 35.8% vs 37.5%). However, both Swin UNETR methods exhibited 4–7% lower AUROC on the near-OOD datasets compared to SMIT-based models, reaffirming that pretraining strategy and encoder architecture substantially influence OOD detection performance within the RF-Deep framework.
5.3 Why RF-Deep outperforms alternatives
(a) Deep features
(b) Radiomics features
Figure 4: t-SNE projected embeddings showing dataset-wise separability of (a) deep features and (b) radiomics features. Results from one representative split (of 100), combining all datasets in a single visualization, are shown for brevity. The shaded blue regions indicate the convex hull of the ID dataset.
5.3.1 RF-Deep extracts features that are better separated than RF-Radiomics
Figure 4 visualizes the t-SNE embeddings (perplexity=30, 1,000 iterations; consistent across random seeds) of deep features (panel a) and radiomics features (panel b) for the ID and OOD datasets collectively. RF-Deep features exhibit better cluster separation between the ID and OOD datasets, with minimal overlap between ID and anatomically distinct far-OOD datasets (KiTS, PancreasCT) and moderate overlap with near-OOD datasets (RSNA PE, MIDRC C19), consistent with their increased detection difficulty. Radiomics features produced poorly separated clusters for the various datasets particularly for the near-OOD datasets with respect to the ID scans, explaining the relatively lower accuracies (Table 1).
We next evaluated whether deep features produced by other pretraining strategies achieved similar separability as the SMIT-pretrained model. As shown in Figure 11, all pretraining methods improved separability compared to handcrafted radiomics features. However, class separation was slightly better using SMIT-based RF-Deep compared to SMIT Lite-based RF-Deep. The differences in class separation were more pronounced between Swin UNETR and Swin UNETR, than among SMIT, SimMIM, and iBOT, indicating the larger impact of pretraining data rather than the pretraining approach itself.
5.3.2 RF-Deep extracts OOD data-agnostic and specific features for improved separability
We used SHAP (SHapley Additive exPlanations) (Lundberg and Lee, 2017) to identify which features drive RF-Deep and RF-Radiomics predictions, revealing how the two approaches differ in their decision-making mechanisms. Deep features (Figure 5) show broad, dataset-specific spread of important dimensions, with some features reused across multiple datasets (Ft 65, Ft 131, Ft 179, Ft 319), suggesting they capture general distributional properties relevant to distinguishing ID from OOD scans. Other features activate selectively for specific distribution shifts (e.g., Ft 5 and Ft 1381 for far-OOD), indicating that RF-Deep adapts its decision boundaries to the characteristics of each OOD scenario rather than relying on a fixed, universal subset of features. Notably, only a small fraction of the 1,488 encoder features appear consistently useful across OOD datasets, underscoring the need for outlier exposure, especially in near-OOD scenarios where subtle feature combinations are required to distinguish concept shifts.
By contrast, radiomics features (Figure 6) emphasized shape-based and texture-based descriptors, specifically the Gray-Level Co-occurrence Matrix (GLCM) and Gray-Level Run-Length Matrix (GLRLM) subsets. While radiomic features lend interpretability compared to deep features, they are insufficient to capture the subtle differences across the datasets to accurately detect OOD.
(a) RSNA PE
(b) MIDRC C19
(c) KiTS
(d) PancreasCT
Figure 5: SHAP feature importance for RF-Deep across OOD datasets using the SMIT-pretrained segmentation model. Different deep features are activated for different distribution shifts.
(a) RSNA PE
(b) MIDRC C19
(c) KiTS
(d) PancreasCT
Figure 6: SHAP feature importance for RF-Radiomics across OOD datasets using segmentations from the SMIT-pretrained model. RF-Radiomics primarily relies on GLCM and GLRLM texture descriptors.
5.3.3 Why logit-based spatial uncertainties are less accurate than RF-Deep?
Figure 7: Representative MaxLogit visualizations for (a) segmentation predictions, (b) spatial heatmaps, and (c) boundary vs. interior distributions with predicted tumor volumes. Violet regions indicate unevaluated voxels due to foreground thresholding.
Table 2: MaxLogit scores across ID and OOD datasets within predicted tumor (positive class) regions, stratified into overall, boundary, and interior voxels.
Dataset Overall Boundary Interior
ID 2.63 $\pm$ 0.80 1.13 $\pm$ 0.35 3.25 $\pm$ 0.92
RSNA PE 1.42 $\pm$ 0.56 0.74 $\pm$ 0.25 1.92 $\pm$ 0.62
MIDRC C19 1.27 $\pm$ 0.70 0.82 $\pm$ 0.41 2.01 $\pm$ 0.73
KiTS 1.09 $\pm$ 0.40 0.85 $\pm$ 0.34 1.57 $\pm$ 0.49
PancreasCT 1.08 $\pm$ 0.60 0.66 $\pm$ 0.32 1.55 $\pm$ 0.75
Table 3: Classifier comparison on the pooled deep features. Point estimates of AUROC ($\uparrow$) and FPR95 ($\downarrow$) are reported over 100 matched-seed runs.
Dataset Linear Probing MLP Classifier Random forest
RSNA PE 92.76 / 28.00 93.01 / 19.31 95.80 / 15.20
MIDRC C19 92.36 / 29.46 92.88 / 27.88 93.30 / 25.50
KiTS 98.92 / 0.110 99.47 / 2.490 100.0 / 0.10
PancreasCT 99.94 / 0.460 99.25 / 0.800 100.0 / 0.00
To understand why logit-based OOD detectors were less accurate than RF-Deep, we analyzed MaxLogit spatial patterns within predicted tumor regions. We extracted boundary regions as a ring of 3 voxel thickness by performing erosion followed by dilation, each using a radius of 1 voxel. The interior contour following the erosion operation constituted the interior regions.
MaxLogit scores of ID and OOD scans show strong overall separation (Mann-Whitney U tests, all p << 0.001, effect sizes ranging from r = -0.75 for MIDRC C19 to r = -0.90 for KiTS) due to higher overall MaxLogit scores for ID scans than individual OOD datasets (Table 2). However, the scores have substantially overlapping ranges with ID (ID: 0.12–4.33, RSNA PE: 0.09–3.67, MIDRC C19: 0.30–2.70).
Concretely, interior voxels showed consistently higher MaxLogit scores than the boundary voxels in both ID and OOD predictions. For ID scans, the interior scores were roughly three times higher than boundary scores, and this ratio persists across all OOD datasets despite their lower absolute scores. Paired, two-sided Wilcoxon signed-rank tests comparing interior and boundary scores within each dataset confirmed the universality of this spatial pattern (p << 0.001 for all datasets, effect sizes r $\approx$ -1.0), indicating that the boundary–interior relationship is preserved regardless of ID or OOD status, further limiting their discriminative power.
Figure 7 provides a visualization of the MaxLogit spatial uncertainties for four representative scans. The density-normalized histograms depict the distribution of MaxLogit values in the interior and boundary regions of the detected lesions. The interior and boundary regions show very good separation for the ID case (row 1) and sufficiently different spatial distribution for the second case (row 2), enabling correct detection of OOD for this case. However, the third case (row 3) presents a challenge wherein the spatial uncertainty distribution closely resembles that of the ID case, resulting in incorrect classification. In contrast, RF-Deep correctly identified both OOD scans, demonstrating its superior ability to capture contextual features beyond local spatial patterns. Of note, both RF-Deep and MaxLogit-based OOD prediction was wrong for the fourth case (row 4), presenting a limiting case for either method.
An additional limitation of logit-based methods is their inherent scale invariance: MaxLogit scores reflect local appearance characteristics independent of lesion size as shown by the distributions for lesions of varying sizes in Figure 7. This is because MaxLogit computation treats voxels identically based on their pre-softmax logit values, ignoring volumetric context that could differentiate true tumors from anatomically implausible detections. This scale invariance poses a critical bottleneck for OOD detection, as many false-positive detections on OOD scans manifest as small, scattered regions with locally high confidence but globally inconsistent spatial distributions. RF-Deep, on the other hand, aggregates features from regions encompassing detections, reducing the impact of tumor size variations.
5.4 Framework generalization
5.4.1 Effect of classifier selection
Table 3 compares different classifiers applied to the same set of encoder features. While linear probing and multilayer perceptron (MLP) classifiers, both implemented with default settings from scikit-learn, provide reasonable separability on the far-OOD datasets, their performance degrades notably on the near-OOD datasets (AUROC $\approx 92--93\%$, FPR95 $\approx 28--30\%$). In contrast, random forests achieve the best performance across all datasets, highlighting the advantage of tree-based non-linear classifiers for capturing complex decision boundaries, particularly when distribution shifts are subtle (Figure 4(a)). The perfect performance on KiTS and PancreasCT (AUROC = 100%, FPR95 $\approx$ 0%) reflects the very easy nature of far-OOD detection when anatomical regions differ substantially from the training distribution.
5.4.2 Generalization to unseen OOD categories
Table 4: Generalization to unseen OOD categories across alternative training strategies on the random forests. Point estimates and 95% bootstrap confidence intervals over 100 matched-seed runs are reported.
Strategy AUROC $\uparrow$ FPR95 % $\downarrow$
RSNA PE MIDRC C19 KiTS PancreasCT RSNA PE MIDRC C19 KiTS PancreasCT
Dataset-specific
95.80
(92.60–98.50)
93.30
(89.90–96.00)
100.00
(99.90–100.00)
100.00
(100.00–100.00)
15.20
(7.10–23.50)
25.50
(16.80–36.20)
0.10
(0.00–1.00)
0.00
(0.00–0.00)
Ensemble
93.70
(91.70–96.50)
92.80
(89.80–95.60)
100.00
(99.80–100.00)
100.00
(100.00–100.00)
19.90
(11.20–29.10)
32.70
(20.40–54.10)
0.00
(0.00–0.00)
0.00
(0.00–0.00)
Unified
93.00
(85.50–97.50)
91.40
(85.40–94.90)
99.80
(99.00–100.00)
99.90
(99.30–100.00)
24.10
(11.20–42.90)
37.30
(20.90–69.40)
0.70
(0.00–5.60)
0.50
(0.00–4.10)
LODO
91.20
(85.50–95.30)
91.00
(84.40–94.80)
99.80
(99.30–100.00)
100.00
(99.70–100.00)
25.70
(17.30–36.70)
43.20
(23.50–54.10)
0.70
(0.00–4.10)
0.20
(0.00–2.00)
LODO+
92.10
(87.50–95.30)
91.80
(84.40–94.80)
99.80
(99.30–100.00)
100.00
(99.70–100.00)
21.70
(15.30–28.70)
34.20
(25.50–48.10)
0.70
(0.00–4.10)
0.20
(0.00–2.00)
We evaluated RF-Deep’s ability to generalize to previously unseen OOD categories using four additional training strategies (Table 4): (i) ensemble, (ii) unified, (iii) leave-one-dataset-out (LODO), and (iv) LODO+. Dataset-specific (DS; baseline) trained a separate RF classifier for each OOD dataset, representing an upper-bound setting where the OOD type was known during development and evaluation. Ensemble trained the same four dataset-specific classifiers but averaged their predictions at test time, assuming no knowledge of the test domain during testing. Unified replaced the per-dataset training with a single RF classifier trained on ID versus all OOD datasets combined. LODO followed the standard OOD paradigm by training on ID plus three OOD datasets and evaluating on the held-out fourth, directly measuring generalization to novel OODs. Finally, LODO+ extended LODO by incorporating randomly sampled non-tumor-centric ROIs to increase anatomical diversity during training.
Performance patterns on near-OOD datasets (RSNA PE and MIDRC C19) revealed important insights on within-anatomy generalization. The LODO strategy showed notable degradation compared to the dataset-specific baseline (RSNA PE: AUROC 91.20 vs 95.80, FPR95 25.70% vs 15.20%; MIDRC C19: AUROC 91.00 vs 93.30, FPR95 43.20% vs. 25.50%). This indicates that training on a single thoracic pathology provides insufficient diversity to generalize to pathologies occurring in the same anatomy. However, LODO+ improved performance over standard LODO (RSNA PE: FPR95 21.70% vs. 25.70%; MIDRC C19: FPR95 34.20% vs. 43.20%), demonstrating that increasing anatomical sampling diversity during training enhanced generalization to unseen near-OOD categories, which is critical for robust OOD detection.
The unified approach, which trained on all thoracic pathologies simultaneously, achieved intermediate performance between LODO and dataset-specific settings (RSNA PE: AUROC 93.00, MIDRC C19: AUROC 91.40), confirming that exposure to multiple within-anatomy pathologies during training improves cross-generalization within the same anatomical region. The ensemble strategy further improved near-OOD performance, approaching dataset-specific performance while maintaining the ability to detect novel OOD categories without prior knowledge of the test domain. On far-OOD datasets (KiTS and PancreasCT), all strategies achieved near-perfect performance (AUROC $\geq$ 99.80, FPR95 $\leq$ 0.70%), with minimal degradation from the dataset-specific baseline.
5.5 Design choices and sensitivity analysis
(a) Number of tumor-anchored ROIs
(b) Crop size of ROIs
(c) Number of training examples
(d) Individual stages of the Swin Transformer
Figure 8: Ablation studies showing the sensitivity of RF-Deep for OOD detection across four datasets. Point estimates and 95% bootstrap confidence intervals over 100 runs with matched seeds are displayed.
To understand the design choices of our proposed OOD detector, we conducted a series of controlled ablations on the number of ROIs, their crop sizes, the extent of labeled (supervised) data, and the encoder stages used for feature extraction. Figure 8(a) shows that OOD detection benefits from richer local context around predicted tumor regions, but only up to a point. With $n=1$, performance is unstable, reflected in wider confidence intervals and elevated FPR95, suggesting that a single tumor-anchored crop may miss informative variation in tumor surroundings. Increasing the number of ROIs from one to four improved performance on near-OOD scans, indicating that additional samples may better disambiguate similarly appearing tumor (ID) from non-cancerous structures. Beyond four ROIs, gains plateau or occasionally regress (e.g., MIDRC C19), likely due to noise injection from ROIs sampling redundant or low-value crops. Far-OOD datasets did not exhibit performance gains with increasing number of ROIs. Therefore, we used $n=4$ to capture meaningful contextual diversity while maintaining computational efficiency.
Figure 8(b) analyzes the effect of tumor-anchored 3D crop size on OOD detection performance. Performance remained relatively stable on the far-OOD datasets across crop sizes, with slight preference for $128\times 128\times 128$. On the near-OOD datasets, both smaller (<128<128) and larger (>128>128) ROIs led to performance degradation, suggesting that the optimal contextual window balances tumor-centered information with surrounding anatomical context.
Figure 8(c) shows that RF-Deep’s OOD detection performance improved with the number of labeled examples, using equal numbers of ID and OOD scans (5/5, 10/10, up to 20/20, corresponding to less than 30% of available data), after which performance plateaued. With only 20 ID and 20 OOD labeled scans, RF-Deep achieved AUROC >> 90 and FPR95 << 30% on average across all OOD datasets. Performance gains were more noticeable on the near-OOD datasets than the far-OOD datasets, indicating that detection on subtle shifts in the same anatomical region is trickier and requires more examples than larger shifts across different anatomical regions. Variance decreases consistently as more samples are added, with the strongest variance reduction observed on near-OOD datasets.
Figure 8(d) shows individual Swin Transformer stage contributions. Far-OOD datasets were relatively insensitive to individual stage selection (achieving AUROC >> 98 even when using features from a single stage), whereas the near-OOD counterparts showed strong sensitivity to early stages, with Stage 1 and Stage 2 features being particularly important (AUROC $\approx$ 88–90 individually). Concatenating features across all stages yielded the best performance, confirming that hierarchical representations are essential for robust OOD detection. This finding aligns with the SHAP analysis (Section 5.3.2), which showed that RF-Deep adaptively leverages features from different abstraction levels depending on the OOD scenario.
6 Discussion and conclusion
The main finding of this work is that effective near-OOD detection requires both discriminative features and learned decision boundaries. Despite using identical deep features, the unsupervised Mahalanobis distance (no OOD scans) performed substantially worse than RF-Deep on the near-OOD datasets, demonstrating that RF-Deep’s outlier exposure adoption and non-linear feature selection are essential for distinguishing pathologies within the same anatomical site. RF-Deep improved over logit-based methods, indicating that over-reliance on the ID model to extract confidence boundaries may be unreliable in OOD scans. Although this issue could be addressed to an extent through confidence calibration, prior studies show limitations of such methods (Larrazabal et al., 2023; Mehrtash et al., 2020). RF-Deep overcomes this limitation by combining SSL-pretrained backbones for segmentation with lightweight random forest classifiers, an approach that is agnostic to the choice of pretraining approach as well as model size, thus providing a plug-and-play approach for leveraging an existing task-specific DL method for OOD detection.
Potential biases and mitigation. To mitigate data-specific biases, we employed: (a) SSL pretraining on 10,432 diverse CT scans encompassing multiple institutions and scanner manufacturers, (b) separate training and testing datasets from different patient datasets providing external validation, and (c) disaggregated evaluation across scanner manufacturer, reconstruction kernels, and contrast agent presence (Figure 3(b)).
Limitations. RF-Deep depends on the segmentation model producing predictions; for the 10% of OOD scans with no foreground output, alternative evaluation strategies (e.g., flagging the absent predictions) would be needed as the datasets expand. While LODO experiments demonstrate generalization to unseen OOD categories (Table 4), near-OOD performance degrades when the held-out category differs substantially from training outliers. Finally, our evaluation was limited to the task of single-label lung cancer segmentation, and generalization to other anatomical sites and multi-label tumors needs to be evaluated.
In summary, we introduced RF-Deep, a post-hoc out-of-distribution (OOD) detection framework for segmentation models that leverages hierarchical encoder features and random forest classifiers to identify concept drifts. Across near-OOD and far-OOD datasets, RF-Deep consistently outperformed logit-based methods and radiomics feature-based classifiers. Future work will explore extension to multi-label segmentations and integration of the absent-prediction signal evaluation for improved robustness.
Broader Impact Statement
This work addresses a critical gap between research-grade AI performance and clinical deployment requirements for medical image segmentation. By providing reliable OOD detection, RF-Deep can help prevent confidently incorrect predictions from reaching clinical workflows, potentially reducing harm to patients from automated systems that fail silently on out-of-distribution inputs. However, users should be aware that no OOD detection method is perfect; the evaluated datasets capture key failure modes, but covering all possible pathologies is impractical. Clinical deployment should include human oversight and validation on institution-specific data distributions before relying on model predictions for patient care decisions.
Appendix A Pretraining strategies descriptions
SMIT (Jiang et al., 2022) consists of masked voxel token prediction, masked patch token distillation, and image token distillation. Masked voxel token prediction is a dense regression task of the intensities within masked patches and is optimized by minimizing the $\ell_{1}$ norm between the predicted and ground-truth patches, akin to SimMIM (Xie et al., 2022). Self-distillation (Tarvainen and Valpola, 2017) was required for the next two tasks and was performed using an exponentially moving average (EMA) teacher model with identical architecture as the student, similar to iBOT (Zhou et al., 2022). Masked patch token distillation aligns the voxel token representations learned by the student with those extracted by the teacher from unmasked sequences, optimized with a temperature-scaled cross-entropy loss. For image token distillation, the cross-entropy between global volume embeddings of the student and teacher was minimized. After pretraining, the teacher was discarded and the student encoder was retained for fine-tuning.
Similarly, Swin UNETR (Tang et al., 2022) employs three pretraining tasks. The primary pretext task is masked volume reconstruction, where a proportion of 3D input patches are randomly masked and the model is trained to regress their voxel intensities from visible context. Unlike SimMIM, which leverages positional encodings to aid masked token prediction, this reconstruction task provided no explicit positional information, forcing the model to learn semantically plausible inpainting from local context. 3D rotation prediction enforces invariance by classifying the degree of rotation applied along the $z$-axis. Finally, contrastive coding encouraged representation learning by maximizing cosine similarity between augmented views of the same volume while minimizing similarity for different volumes. Unlike SMIT, which requires an EMA teacher, Swin UNETR used only a student model. The pretrained encoder was directly retained for fine-tuning.
Appendix B Logit-based approaches formulae
The tumor segmentation problem can be broken down into a per-voxel binary classification problem, in which we mean-aggregate the scores of the positive class to summarize the scan. In-distribution (ID) scans tended to yield more confident outputs, whereas out-of-distribution (OOD) scans produced less confident or spurious predictions. Following the convention of treating OOD as the positive class (Hendrycks and Gimpel, 2017; Choi et al., 2024), we negated each score so that higher values correspond to greater likelihood of being OOD. Let $f^{i}(x)$ denote the logit for class $i\in\{0,1\}$ at voxel $x$, and define the softmax score as
$S_{\text{softmax}}^{i}(x)\;=\;\frac{\exp\!\big(f^{i}(x)\big)}{\sum_{j=0}^{1}\exp\!\big(f^{j}(x)\big)}.$
We denote the set of voxels predicted as positive (tumor) by
$V_{+}\;=\;\bigl\{(h,w,d)\,\big|\,\hat{y}_{h,w,d}=1\bigr\},$
where $\hat{y}_{h,w,d}$ is the class prediction at voxel $(h,w,d)$. Each logit-based score aggregates over $V_{+}$:
  • MaxSoftmax (Hendrycks and Gimpel, 2017):
    $S\;=\;-\,\frac{1}{|V_{+}|}\sum_{(h,w,d)\in V_{+}}S_{\text{softmax}}^{1}\!\left(x_{h,w,d}\right).$
  • MaxLogit (Hendrycks et al., 2022):
    $S\;=\;-\,\frac{1}{|V_{+}|}\sum_{(h,w,d)\in V_{+}}f^{1}\!\left(x_{h,w,d}\right).$
  • Energy (Liu et al., 2020):
    $S\;=\;-\,\frac{1}{|V_{+}|}\sum_{(h,w,d)\in V_{+}}\log\!\Bigg(\sum_{i=0}^{1}\exp\!\big(f^{i}(x_{h,w,d})\big)\Bigg).$
Appendix C Additional results
This section provides additional quantitative and qualitative results that support the findings presented in the paper.
  • Figure 9 presents the disaggregated analysis of segmentation performance across imaging variations, discussed briefly in Section 5.1 (Figure 3).
  • Table 5 lists the mapping from feature IDs used across Sections 5.2 and 5.3 to standardized IBSI feature names (retaining only those post recursive feature elimination).
  • Figure 10 supports the discussion in Section 5.2 on MD-Deep, demonstrating that the standard formulation outperforms all variants.
  • Table 6 reports the complete OOD detection performance of RF-Deep across all six evaluated backbones, briefly discussed in Section 5.2.
  • Table 7 summarizes OOD performance for all methods (five additional pretraining strategies on six methods), reinforcing the findings in Section 5.2.
  • Figure 11 shows t-SNE embeddings for all alternative pretraining strategies, complementing Figure 4(a) and discussions in Sections 5.2 and 5.3.1.
Figure 9: Segmentation performance across backbones under imaging variations (scanner type, contrast, reconstruction kernel) on the ID dataset. DSC and HD95 are reported.
Table 5: Radiomics features definitions.
ID Feature ID Feature
General features
feat_5feat\_5 maxIntensityInterpReseg feat_6feat\_6 minIntensityInterpReseg
Shape features
shape_7shape\_7 max2dDiameterSagittalPlane shape_11shape\_11 volume
shape_12shape\_12 filledVolume shape_18shape\_18 surfToVolRatio
First-order features
first_1first\_1 min first_2first\_2 max
first_4first\_4 range first_10first\_10 entropy
first_12first\_12 energy first_13first\_13 totalEnergy
Gray-Level Co-occurrence Matrix (GLCM) features
glcm_8glcm\_8 jointEntropy_3D_StdDev glcm_18glcm\_18 jointAvg_3D_StdDev
glcm_36glcm\_36 sumEntropy_3D_avg glcm_37glcm\_37 sumEntropy_3D_Median
glcm_39glcm\_39 sumEntropy_3D_Min glcm_43glcm\_43 contrast_3D_StdDev
glcm_44glcm\_44 contrast_3D_Min glcm_47glcm\_47 invDiffMom_3D_Median
glcm_48glcm\_48 invDiffMom_3D_StdDev glcm_58glcm\_58 invDiff_3D_StdDev
glcm_59glcm\_59 invDiff_3D_Min glcm_60glcm\_60 invDiff_3D_Max
glcm_61glcm\_61 invDiffNorm_3D_avg glcm_63glcm\_63 invDiffNorm_3D_StdDev
glcm_68glcm\_68 invVar_3D_StdDev glcm_76glcm\_76 diffEntropy_3D_avg
glcm_77glcm\_77 diffEntropy_3D_Median glcm_78glcm\_78 diffEntropy_3D_StdDev
glcm_79glcm\_79 diffEntropy_3D_Min glcm_80glcm\_80 diffEntropy_3D_Max
glcm_81glcm\_81 diffVar_3D_avg glcm_82glcm\_82 diffVar_3D_Median
glcm_84glcm\_84 diffVar_3D_Min glcm_89glcm\_89 diffAvg_3D_Min
glcm_92glcm\_92 corr_3D_Median glcm_93glcm\_93 corr_3D_StdDev
glcm_95glcm\_95 corr_3D_Max glcm_98glcm\_98 clustTendency_3D_StdDev
Gray-Level Run-Length Matrix (GLRLM) features
glrlm_28glrlm\_28 runLengthNonUniformityNorm_3D_StdDev glrlm_68glrlm\_68 grayLevelVariance_3D_StdDev
glrlm_73glrlm\_73 runLengthVariance_3D_StdDev glrlm_78glrlm\_78 runEntropy_3D_StdDev
Gray-Level Size Zone Matrix (GLSZM) features
glszm_3glszm\_3 grayLevelNonUniformity_3D
Gray-Tone Difference Matrix (GTDM) features
gtdm_5gtdm\_5 strength_3D
Table 6: RF-Deep OOD detection performance across different backbone encoders. AUROC ($\uparrow$) and FPR95 ($\downarrow$) with 95% bootstrap confidence intervals are reported over 100 matched-seed runs.
Backbone RSNA PE MIDRC C19 KiTS PancreasCT
AUROC FPR95 % AUROC FPR95 % AUROC FPR95 % AUROC FPR95 %
SimMIM
93.70
(90.40–96.10)
17.60
(10.20–28.10)
93.20
(90.10–96.10)
26.50
(18.30–38.20)
99.90
(99.80–100.00)
0.10
(0.00–1.00)
99.90
(99.80–100.00)
0.60
(0.00–2.00)
iBOT
93.60
(91.70–96.80)
22.60
(12.20–42.40)
92.70
(89.30–96.30)
24.60
(14.70–36.70)
99.80
(99.30–100.00)
1.10
(0.00–4.10)
99.80
(98.90–100.00)
1.10
(0.00–7.70)
SMIT Lite
94.10
(90.10–96.10)
20.40
(11.20–34.30)
90.40
(86.70–94.10)
30.50
(20.40–41.80)
99.60
(99.30–99.90)
1.80
(0.50–3.10)
99.30
(98.80–99.90)
1.80
(0.00–4.10)
Swin UNETR
89.60
(85.10–92.90)
34.10
(20.40–51.10)
88.90
(84.70–92.20)
37.50
(25.00–52.60)
99.70
(99.20–99.90)
2.40
(0.00–5.10)
99.60
(99.00–99.90)
2.90
(0.00–6.10)
Swin UNETR
90.40
(84.70–94.10)
41.50
(25.00–58.20)
88.60
(83.50–92.30)
35.80
(24.50–50.00)
99.40
(98.60–99.90)
2.60
(0.00–6.10)
98.70
(97.00–99.80)
7.10
(1.00–18.90)
SMIT
95.80
(92.60–98.50)
15.20
(7.10–23.50)
93.30
(89.90–96.00)
25.50
(16.80–36.20)
100.00
(99.90–100.00)
0.10
(0.00–1.00)
100.00
(100.00–100.00)
0.00
(0.00–0.00)
Figure 10: Ablation of Mahalanobis distance-based variants for OOD detection using AUROC over 100 matched-seed runs.
Table 7: RF-Deep OOD detection performance across different pretrained backbone encoders. AUROC ($\uparrow$) and FPR95 ($\downarrow$) with 95% bootstrap confidence intervals are reported over 100 matched-seed runs.
Model Method RSNA PE MIDRC C19 KiTS PancreasCT
AUROC FPR95 % AUROC FPR95 % AUROC FPR95 % AUROC FPR95 %
SimMIM MaxSoftmax 86.89 45.56 84.58 54.92 94.19 21.59 94.30 24.04
MaxLogit 89.39 44.04 87.55 45.65 95.99 15.32 96.28 14.82
Energy 89.70 44.64 87.75 45.24 96.15 14.92 96.39 13.84
RF-Radiomics 87.20 48.90 89.80 34.90 97.00 15.80 96.60 17.90
MD-Deep 88.60 41.32 86.85 51.72 97.66 3.40 97.84 8.30
RF-Deep 93.70 17.60 93.20 26.50 99.90 0.10 99.90 0.60
iBOT MaxSoftmax 83.56 48.25 80.97 52.10 90.45 32.22 88.98 36.23
MaxLogit 87.31 38.55 85.24 39.33 93.82 23.45 92.29 27.17
Energy 87.90 38.62 85.84 39.61 94.49 22.02 92.96 25.17
RF-Radiomics 89.30 50.90 90.50 29.70 96.80 14.70 96.80 15.20
MD-Deep 87.83 44.87 86.12 52.89 98.63 6.62 96.16 17.27
RF-Deep 93.60 22.60 92.70 24.60 99.80 1.10 99.80 1.10
SMIT Lite MaxSoftmax 77.96 56.05 74.44 60.96 80.77 48.42 87.35 41.45
MaxLogit 78.52 53.32 79.03 51.95 79.46 45.97 85.42 41.33
Energy 78.16 53.82 79.37 52.22 78.74 45.59 83.90 40.53
RF-Radiomics 88.90 38.90 90.70 31.10 97.10 13.80 97.40 12.80
MD-Deep 85.94 38.85 78.51 56.77 93.51 25.57 92.26 25.06
RF-Deep 94.10 20.40 90.40 30.50 99.60 1.80 99.30 1.80
Swin UNETR MaxSoftmax 80.30 52.19 79.65 62.01 84.40 46.38 88.97 46.98
MaxLogit 83.88 42.62 83.85 52.12 88.99 34.86 91.57 35.45
Energy 84.16 40.77 84.30 50.44 89.41 31.86 92.17 33.91
RF-Radiomics 87.70 37.40 89.20 33.40 98.10 9.80 98.10 13.70
MD-Deep 77.88 62.47 74.14 64.61 92.61 25.50 89.54 29.67
RF-Deep 89.60 34.10 88.90 37.50 99.70 2.40 99.60 2.90
Swin UNETR MaxSoftmax 75.92 53.25 77.83 55.91 78.89 53.17 93.12 23.96
MaxLogit 77.87 43.65 80.41 49.93 83.33 39.30 92.37 18.44
Energy 78.02 42.46 80.78 49.48 83.87 38.24 92.17 18.98
MD-Deep 74.03 68.07 70.50 71.02 84.56 49.44 79.45 63.34
RF-Radiomics 86.50 40.00 89.00 37.20 97.40 11.30 97.80 11.70
RF-Deep 90.40 41.50 88.60 35.80 99.40 2.60 98.70 7.10
(a) SimMIM
(b) iBOT
(c) SMIT Lite
(d) Swin UNETR
(e) Swin UNETR
Figure 11: t-SNE projected embeddings showing dataset-wise separability of (a) deep features and (b) radiomics features. Results from one representative split (of 100), combining all datasets in a single visualization, are shown for brevity. The shaded blue regions indicate the convex hull of the ID dataset.